### v1

# KG-QAGen: A Knowledge-Graph-Based Framework for Systematic Question Generation and Long-Context LLM Evaluation

Abstract
The increasing context length of modern language models has created a need for evaluating their ability to retrieve and process information across extensive documents. While existing benchmarks test long-context capabilities, they often lack a structured way to systematically vary question complexity. We introduce KG-QAGen (Knowledge-Graph-based Question-Answer Generation), a framework that (1) extracts QA pairs at multiple complexity levels (2) by leveraging structured representations of financial agreements (3) along three key dimensions – multi-hop retrieval, set operations, and answer plurality – enabling fine-grained assessment of model performance across controlled difficulty levels. Using this framework, we construct a dataset of 20,139 QA pairs (the largest number among the long-context benchmarks) and open-source a part of it. We evaluate 13 proprietary and open-source LLMs and observe that even the best-performing models are struggling with set-based comparisons and multi-hop logical inference. Our analysis reveals systematic failure modes tied to semantic misinterpretation and inability to handle implicit relations.
1 Introduction
The increasing context length of recent language models has led to growing interest in evaluating their capabilities. Researchers have studied challenges like the Lost in the Middle problem (Liu et al., 2024b), examined the limiting factors of scaling models to long context (such as the constraints imposed by RoPE’s base value (Men et al., 2024)), and developed techniques to extend context length efficiently (Hooper et al., 2024; Ding et al., 2024). Others have focused on improving inference efficiency (Jiang et al., 2024b; Tang et al., 2024) or enhancing long-context utilization and mitigating information loss (Zhang et al., 2024d; Lin et al., 2024; Zhang et al., 2024c).
Numerous benchmarks have been developed to evaluate long-context LLMs, covering a variety of tasks and capabilities. Some benchmarks support multi-lingual evaluation (Agrawal et al., 2024; Bai et al., 2024), include tasks such as segment sorting (Wang et al., 2024a; Dong et al., 2024), numerical reasoning (Zhao et al., 2024c; Reddy et al., 2024), timeline reordering (Zhang et al., 2024a). However, certain aspects of understanding long-context remain underexplored in existing benchmarks.
Figure 1: Overview of KG-QAGen. Credit agreements are annotated to identify entities and their relationships, forming a knowledge graph representation. This graph is then used to systematically extract multi-level QA pairs, which serve as the basis for benchmarking long-context LLMs.
Real-World Documents
Understanding large, real-world documents presents a unique challenge due to complex entity relationships, co-reference variations, and cross-referenced information scattered across sections (Zaporojets et al., 2022). This complexity makes credit agreements a precise fit for long-context evaluation, as they contain networks of entities referenced in varied ways across clauses (Section 3.1) and introduce numerous challenging edge cases (Appendix A).
Question Complexity Levels
Long-context benchmarks can offer deeper insight by distinguishing question difficulty levels, enabling analysis of how performance degrades with increasing logical inference depth – whether due to retrieval limits, context length, or multi-step inference. While some benchmarks contain tasks of varying difficulty (Gupta et al., 2024; Zhao et al., 2024c; Zhu et al., 2024), none explicitly categorize question complexity levels. Moreover, our benchmark not only defines these levels but also provides a systematic approach to creating granular questions, leveraging our annotations to deterministically construct QA pairs with clearly defined complexity directions.
We introduce KG-QAGen (Knowledge-Graph-based Question-Answer Generation), a framework for structured extraction of QA pairs from manually annotated documents via knowledge-graph representations. Our approach structures questions along three dimensions: the number of hops, the use of set operations, and whether the answer is singular or plural. Using a knowledge-graph-inspired representation, we create diverse questions and extract answers efficiently with SPARQL queries. This structured approach ensures scalability, flexibility, and controlled dataset creation, enabling a more precise evaluation of long-context LLMs.
We apply KG-QAGen to 170 credit agreements, resulting in KG-QAGen-D, a dataset of 20,139 QA pairs categorized by three complexity levels. We use it to evaluate 13 long-context LLMs and find that even on simple factual queries, some models fail to extract basic information; performance further drops with multi-hop paths and compositional questions involving set operations (Section 4.3). Our analysis shows that set-based comparisons are especially challenging, with high rates of “Not Found" responses and semantic misinterpretation (Section 5.2).
Our contributions are as follows: (1) We present a framework grounded in knowledge graphs, enabling systematic extraction of QA pairs at varying complexity levels. (2) We open-source a portion of the KG-QAGen-D to encourage future work (see Appendix F for details). (3) We release our inference pipeline and benchmarking codes to facilitate reproducible evaluation of long-context models.
2 Related Work
Table 1: Comparison of KG-QAGen-D with existing long-context LLM benchmarks. Our benchmark stands out by evaluating question complexity on a granular level, incorporating real-world SEC credit agreements, and supporting multi-hop and context-aware retrieval.
Real-world Multi-hop Context-aware Ques. complexity
Dataset documents retrieval retrieval levels Document Type # Samples
mLongRR (Agrawal et al., 2024) partially BBC News unspecified
L-Eval (An et al., 2024) partially Earnings Calls + Misc 2,008
CLongEval (Qiu et al., 2024) partially THUnews + Misc 7,267
NOCHA (Karpinska et al., 2024) Fictional Books 2,002
Marathon (Zhang et al., 2024a) Prior Datasets 1,530
[…]
M4LE (Kwan et al., 2024) partially News + Misc unspecified
LooGLE (Li et al., 2024) partially arXiv + Misc 6,448
BAMBOO (Dong et al., 2024) partially ACL Anthology + Misc 1,502
LongBench (Bai et al., 2024) Previous Datasets 4,750
MMLongBench-DOC (Ma et al., 2024) partially arXiv + Misc 1,082
Loong (Wang et al., 2024b) SEC + Misc 1,600
FinDVer (Zhao et al., 2024b) SEC 10-K & 10-Q 2,100
DocFinQA (Reddy et al., 2024) SEC 10-K & 10-Q 7,437
(Gupta et al., 2024) partially Financial News 560
FanOutQA (Zhu et al., 2024) partially Wikipedia 7,305
DocMath-Eval (Zhao et al., 2024c) partially SEC 10-K & 10-Q 4,000
KG-QAGen-D (ours) SEC Credit Agreements 20,139
Structured retrieval from long-context documents remains a key challenge, especially when handling complex entities and their relationships (Lin et al., 2025; Guo et al., 2024; Saad-Falcon et al., 2023). Multi-hop retrieval and context-aware models are particularly relevant for real-world documents.
Multi-Hop Retrieval
Multi-hop retrieval involves retrieving multiple interconnected pieces of evidence across different sections of a document or multiple documents to answer a question. This capability is crucial for handling complex queries requiring multi-step retrieval, such as tracing relationships between entities in legal documents or extracting financial conditions scattered across an agreement (Wang et al., 2025; Sarmah et al., 2024). Some existing benchmarks (An et al., 2024; Wang et al., 2024a) focus on long-context understanding without incorporating explicit multi-hop retrieval challenges.
Context-Aware Retrieval
Context-aware retrieval refers to tasks where models retrieve relevant information to answer questions based on document content rather than artificial information placement. Some existing benchmarks focus on the Needle in a Haystack problem (Agrawal et al., 2024), where small pieces of information are inserted into long texts and then retrieved. This type of retrieval does not reflect real-world scenarios where understanding the document’s structure and semantics is crucial (Munikoti et al., 2023; Lai et al., 2024).
3 The KG-QAGen Benchmark
KG-QAGen systematically extracts structured question-answer pairs across varying complexity levels for long-context evaluation. Using credit agreements (Section 3.1) annotated with a structured label schema (Section 3.2, Appendix A), we construct a knowledge graph (Section 3.3) to enable systematic query construction. Questions are designed along three dimensions – multi-hop retrieval, set operations, and answer plurality (Section 3.4) – ensuring scalable, diverse, and automated question extraction (Section 3.5). The dataset balances complexity and document coverage, making it a robust benchmark for assessing long-context retrieval.
3.1 Credit Agreements
We sampled 17 credit agreements per year from 2013 to 2022 from SEC EDGAR, totaling 170 documents. Table 2 summarizes their length by token and sentence count. Their suitability for long-context LLM benchmarking is discussed below, with supporting examples in Appendix A.
Rich in Entities and Relations
Credit agreements contain multiple organizations, individuals, roles, and locations, with entities appearing in varied forms throughout the document. Some roles have sub-roles (e.g., a lender can be a Swing-Line Lender or Lead Lender). Extracting information requires cross-referencing sections, making retrieval a challenging task.
Ambiguous Co-reference
Entities are often referenced in multiple ways; for example, “Bank of America, N.A.” may also appear as “Bank of America”. Extracting the exact full name can be challenging, as abbreviations or partial mentions may appear throughout the document.
Structured but not Standardized
While they follow a general structure (e.g., loan terms, covenants, defaults), variations in phrasing and formatting mean that the same type of information can appear in different sections or be expressed in diverse ways, preventing models from relying on fixed templates and requiring deeper understanding of the document’s content.
Real-World Significance
Credit agreements outline multi-million-dollar transactions and define the rights and responsibilities of involved parties. Accurate analysis of these documents is essential for financial institutions, legal teams, and regulators to assess risk, ensure compliance, and avoid costly misunderstandings. This makes them a highly relevant and impactful target for developing and evaluating QA systems.
3.2 Annotations
Table 2: Statistics of the source documents and their structured annotations used to construct KG-QAGen-D. The table summarizes token and sentence counts, as well as the distribution of annotated entities and relations per document. The low minimum values are due to an edge case discussed in Appendix A.3.
Stats per doc (total 170 docs) Min Avg Max
Documents
Tokens 32,562 83,694 233,207
Sentences 470 1,066 3,541
Entities
Org Name 2 27.71 390
Org Role 1 13.88 52
Org Sub-Role 0 4.26 35
Person Name 2 12.55 88
Person Position 2 13.20 88
Location 1 5.62 168
Location Type 1 0.43 7
Relations
Org Name $\rightarrow$ Org Role 1 21.17 79
Org Role $\rightarrow$ Org Sub-Role 0 3.74 32
Org Name $\rightarrow$ Person Name 2 0.97 14
Person Name $\rightarrow$ Person Position 2 3.04 54
Org Name $\rightarrow$ Location 1 5.64 178
Location $\rightarrow$ Location Type 1 0.45 18
Our annotation schema captures key entities and relationships in credit agreements, defining seven label types covering persons, organizations, roles, and locations to represent their structure. Appendix A outlines label definitions, placement rules, and relationship guidelines, while Table 2 illustrates entity relationships and their frequency across documents.
We annotated the extracted credit agreements using Label-Studio [^1] https://github.com/HumanSignal/label-studio, supporting entity linking and relation extraction for .html files. The annotation was carried out by annotators with backgrounds in Computer Science or Mathematics, under the direct supervision of a Master of Finance and a PhD student advised by a Chair Professor of Finance, also a co-author. The annotation followed a structured, iterative approach to ensure consistency and resolve edge cases. The team first established label definitions and completed an initial annotation round. The results were reviewed to identify and discuss edge cases, which led to refinements in the annotation guidelines. To enhance reliability, two additional rounds of annotation cleaning were conducted, with annotators reviewing documents they had not previously worked on.
3.3 Knowledge Graph Representation
Figure 2: Knowledge graph representation of annotated credit agreements. The ontology layer defines high-level entity types and their sub-classes (e.g., Agent, Lender), while the data layer contains document-specific instances connected via labeled relations (e.g., hasEmployee, isLocationOf). This abstraction supports systematic QA extraction across documents by removing company-specific constraints. All entity names in this figure, including individuals, organizations, and addresses, are fictional and used for illustrative purposes only.
To enable systematic and scalable QA extraction, we transform annotated credit agreements into a knowledge graph representation. This graph encodes entities of people, organizations, roles, and locations, along with their relationships, in a structured and queryable form. Figure 2 illustrates a graph which captures that Jane Smith holds the position of Vice President at ABCDE Inc. – an organization that serves as a Lender as well as an Administrative and Documentation Agent, and has its Headquarters at 10 Peachtree.... By abstracting away document-specific phrasing while preserving these logical links, the graph supports consistent interpretation across agreements and forms the basis for creating questions of varying complexity through structured queries.
We construct the knowledge graph using the Resource Description Framework [^2] https://www.w3.org/TR/rdf11-concepts/ (RDF), which represents data as subject-predicate-object triples, forming a directed labeled graph for knowledge representation. Stored in Turtle (*.ttl) format for readability, this structure ensures scalability, and efficient querying across documents using SPARQL [^3] https://www.w3.org/TR/sparql11-query/. See Appendix B for knowledge graph construction details.
3.4 Question Templates and Complexity Dimensions
We begin by identifying all possible 1-hop QA templates from the knowledge graph, illustrated in Figure 2: questions formed from direct relationships between two entities, such as “What is the position of [Person Name]?" or “Who is the representative of [Org Name]?". These base templates are constrained to yield singular answers, forming the simplest form of retrieval. It should be noted that not all templates are universally applicable – for instance, the latter question is only valid when the document specifies exactly one representative for the given organization. We then systematically expand these templates along three dimensions.
Number of Hops
Increasing retrieval depth via intermediate nodes, e.g., “Who is the [Person Position] of the organization located at [Location]?".
Set Operations
Introducing comparisons across entities using set logic, e.g., “What are the positions held by [Person Name A] but not [Person Name B] or [Person Name C]?".
Plurality
Moving from expecting a singular answer to allowing multiple valid answers, e.g., “What companies are the [(Org Sub-Role +) Org Role] in the agreement?" if multiple roles are held.
Table 3: This table defines the question templates used to construct KG-QAGen-D, structured along three dimensions: plurality (P), number of hops (H), and set operations (#SO). It includes example templates, corresponding knowledge graph query paths, and logical operations involved. These dimensions are used to compute the overall complexity level for each QA pair.
P H #SO Example Template Hop Path Set Operation
0 1 0 What is the position of [Person Name]? Person→Position None
1 1 0 What are the positions of [Person Name]? Person→Position None
0 1 1 What position is held by both [Person A] and [Person B]? Person→Position $A\cap B$
0 1 2 What position does [Person A] hold that [Person B] doesn’t? Person→Position $A\cap\neg B$
0 1 3 What position is held by [Person A] but not [Person B] or [Person C]? Person→Position $A\cap\neg B\cap\neg C$
1 1 1 What are the positions held by both [Person A] and [Person B]? Person→Position $A\cap B$
1 1 2 What are the positions held by [Person A] but not [Person B]? Person→Position $A\cap\neg B$
1 1 3 What are the positions held by [Person A] but not [Person B] or [Person C]? Person→Position $A\cap\neg B\cap\neg C$
0 2 0 Who is the [Position] of [Org]? Org→Pos→Per None
1 2 0 What are the roles of [Org] in the agreement where [Person] is employed? Per→Org→Role None
0 2 1 Who is the [Position A] and [Position B] of [Org]? Org→Pos→Per $A\cap B$
1 2 1 Who are both [Position A]s and [Position B]s of [Org]? Org→Pos→Per $A\cap B$
0 3 0 Who is the [Position] of the company associated with [Location]? Loc→Or→Ps→Pe None
1 3 0 Who are the [Position]s of the company associated with [Location]? Loc→Or→Ps→Pe None
0 3 1 Who is both [Position A] and [Position B] of the company associated with [Location]? Loc→Or→Ps→Pe $A\cap B$
0 3 2 Who is [Position A] but not [Position B] of the company associated with [Location]? Loc→Or→Ps→Pe $A\cap\neg B$
0 3 3 Who is [Position A] and [Position B] but not [Position C] of the company associated with [Location]? Loc→Or→Ps→Pe $A\cap B\cap\neg C$
1 3 1 Who are both [Position A] and [Position B] of the company associated with [Location]? Loc→Or→Ps→Pe $A\cap B$
1 3 2 Who are [Position A]s but not [Position B]s of the company associated with [Location]? Loc→Or→Ps→Pe $A\cap\neg B$
3.5 Extraction of QA Pairs from Templates
Table 3 illustrates how questions are systematically constructed by incrementally increasing complexity across three dimensions introduced in Section 3.4: number of hops (H), plurality (P), and number of set operations (#SO). Each row in the table corresponds to a specific question template, with associated values for P, H, and #SO, the query path in the knowledge graph, and the set operation logic used.
Question instances are constructed by dynamically substituting entities into these templates using SPARQL queries over the knowledge graph described in Section 3.3. This ensures precise and scalable QA pair creation, with answer labels grounded in annotated relationships. While our framework supports arbitrarily complex combinations, we observed that highly compositional templates often become inapplicable to real documents. For instance, there are no individuals in an agreement who hold three distinct positions or no organization with the combination of roles required to fulfill a complex query. Thus, we limit template depth since such complex conditions are typically not present in the source documents. All the created templates can be found in Appendix C.
To support evaluation, we define a composite question complexity level $L\!=\!H\!+\!P\!+\!\#SO$, and group QA pairs into three categories. Easy ($L\!=\!1$): simple, single-hop, singular-answer queries. Medium ($42\!\leqslant\!L\!\leqslant\!$ 4): questions with moderate complexity via hops, set operations, or plurality. Hard ($L\!=\!5$): questions with the most complex combinations of all three dimensions.
The resulting dataset, KG-QAGen-D, comprises 20,139 QA pairs from 170 credit agreements, surpassing existing long-context QA benchmarks in size (Table 1). This scale is achieved through efficient question extraction from structured annotations using reusable templates.
4 Evaluation
In this section, we describe the setup for our experiments. We evaluate multiple long-context LLMs (listed in Section 4.1) across different question complexity levels, as defined in Section 3.4. Performance is measured using four metrics detailed in Section D.2. In addition, we conduct human evaluation of LLM responses on a small sample of QA pairs (see Section D.4). Our prompting strategy is outlined in Section D.1. The results of our level-based evaluation are presented in Table 4, with the analysis provided in Section 5.
4.1 Baselines
We evaluate 13 proprietary and open-source LLMs on KG-QAGen-D benchmark.
Proprietary LLMs:
GPT-4o & GPT-4o-Mini (Achiam et al., 2023; Hurst et al., 2024), Gemini-1.5-Pro & Gemini-1.5-Flash (Team et al., 2023, 2024a), and Gemini-2.0-Flash & Gemini-2.0-Flash-Lite (Team et al., 2024a).
Open-Source LLMs:
DeepSeek-V3 (Liu et al., 2024a), Llama-4-Maverick-17B-128E-Instruct-FP8 (AI, 2024a), Llama-4-Scout-17B-16E-Instruct (AI, 2024b), Llama-3.1-405B-Instruct-Turbo & Llama-3.1-70B-Instruct & Llama-3.1-8B-Instruct (Dubey et al., 2024), and Qwen2-72B-Instruct (Yang et al., 2024).
Given that our documents average 80k tokens, we restrict our evaluation to models supporting a context length of at least 128k tokens. Nonetheless, even long-context models require a carefully designed prompting strategy to extract relevant answers effectively (see Section D.1). We made several attempts to benchmark the recently introduced Gemini-2.5-Pro and Gemini-2.5-Flash (Google, 2025) (available in preview mode), but they consistently returned empty responses regardless of configuration. We also attempted to evaluate Gemma-2-8k (Team et al., 2024b), Mixtral-Small-32k, and Mixtral-8×22-65k (Jiang et al., 2024a), but these models frequently returned “Not found", making systematic evaluation infeasible.
For proprietary models and those accessed via Together.AI [^4] https://www.together.ai/, we utilized the LangChain [^5] https://www.langchain.com/ framework, while open-source models were employed using vLLM (Kwon et al., 2023) for efficient inference. To ensure reproducibility, we set the temperature to 0.0.
4.2 Experiment Setup
Our evaluation setup is designed to handle long financial documents and large batches of questions in a way that fits the context length constraints of modern language models. We evaluate models by prompting them with a document (or a chunk, if needed) and a batch of 50 associated questions, then measuring the quality of their answers.
As seen in Figure 3, if a document exceeds the 128K-token context limit, we split it into manageable chunks. For each chunk or full document (if it fits), we prompt the model with a batch of 50 questions associated with that segment. This design ensures consistent and scalable evaluation across documents of varying lengths. See Appendix D.1 for further details.
We use four evaluation metrics: F1 score, Edit Distance, Cosine Similarity, and LLM-as-a-Judge. Among these, F1 and LLM-as-a-Judge are highlighted in Table 4 for their complementary strengths – F1 quantifies exact token-level overlap, while LLM-as-a-Judge captures semantic correctness with human-aligned scoring. The details are given in Appendix D.2. F1 and LLM-as-a-Judge also show the highest correlation with human evaluation metrics (see Appendix D.4).
Figure 3: This figure illustrates the multi-stage process used to evaluate long-context LLMs: documents are split into chunks (if exceeding 128K tokens), paired with question batches ($\leqslant$50), and sequentially fed to the model. Model responses are collected per chunk and per question batch to ensure evaluation scalability across large, real-world documents.
4.3 Main Results
The results in Table 4 highlight a significant decline in models performance across all metrics as question complexity increases. Llama-3.1-Instruct-Turbo (405B) stands out in the Hard category, achieving an F1 score of 0.3709, significantly higher than all other models, despite not leading in Easy or Medium tasks, which could indicate stronger generalization to difficult tasks. However, it was the only model that failed to follow the requested JSON output format, requiring us to develop a separate parser. This suggests that as models become larger and more general, their ability to follow specific instructions diminishes.
Gemini-2.0-Flash (1000K) and Gemini-1.5-Flash (1000K) dominate in Easy and Medium queries but struggle on Hard, indicating a limitation in handling complex retrieval. Among open-source models, DeepSeek-V3 (671B) performs consistently but lacks dominance, while Llama-3.1-Instruct (70B) competes well on Medium, surpassing some larger proprietary models. Qwen2-Instruct (72B) underperforms across all categories, suggesting size alone does not determine effectiveness.
Table 4: This table presents the performance of 13 LLMs, evaluated across Easy, Medium, and Hard question categories. Metrics include F1 Score and LLM-as-a-Judge rating, capturing both token-level accuracy and semantic correctness. The results reveal a consistent decline in performance as question complexity increases, with notable model-specific strengths and weaknesses. denotes the models evaluated on a smaller subset due to cost constraints.
Easy Medium Hard
Model Size Context F1 LLM-as- F1 LLM-as- F1 LLM-as-
Length Score $\uparrow$ a-Judge $\uparrow$ Score $\uparrow$ a-Judge $\uparrow$ Score $\uparrow$ a-Judge $\uparrow$
Proprietary LLMs
Gemini-2.0-Flash 1M 0.6748 4.0000 0.3294 2.4820 0.1172 2.2838
Gemini-2.0-Flash-Lite 1M 0.5908 3.4172 0.3883 2.6497 0.1943 1.5135
Gemini-1.5-Pro 2M 0.5410 3.2448 0.4577 2.8436 0.1791 1.6986
Gemini-1.5-Flash 1M 0.5103 3.1207 0.3892 2.5130 0.1400 1.5411
GPT-4o 128K 0.5272 3.2034 0.3783 2.5451 0.1178 1.4795
GPT-4o-Mini 128K 0.4723 2.9690 0.2732 2.0681 0.0274 1.1370
Open-source LLMs
DeepSeek-V3 671B 128K 0.6218 3.5668 0.4796 3.0210 0.0856 1.4795
Llama-4-Maverick-Instruct 400B 1M 0.5552 3.4482 0.3848 2.6976 0.2031 3.1351
Llama-4-Scout-Instruct 109B 1M 0.4993 3.2138 0.2215 1.8772 0.0472 1.5270
Llama-3.1-Instruct-Turbo 405B 128K 0.5472 3.4793 0.4204 2.6377 0.3651 2.7297
Llama-3.1-Instruct 70B 128K 0.4955 3.1373 0.4017 2.6722 0.1639 2.3425
Llama-3.1-Instruct 8B 128K 0.3891 2.6996 0.2962 2.2332 0.1169 1.4315
Qwen2-Instruct 72B 128K 0.3886 3.3884 0.3100 2.9078 0.0762 1.3288
5 Analysis
5.1 Overall Error Trends
Figure 4: Trends in error types across complexity levels. The template level is given by $P+H+\#SO$.
As question complexity increases, the LLM’s ability to retrieve and generate correct responses degrades markedly. With increase in number of hops or set operations, the percentage distribution of each error type increases (see Figure 4 for more details).
  • Not Found Responses. Increases from 15.00% at Level 1 to 77.39%, reflecting increasing retrieval difficulty.
  • Low F1 Scores. Rises from 32.55% to 84.25%, indicating less accurate extraction.
  • Low Cosine Similarity. Jumps from 38.03% to 92.47%, showing semantic drift.
  • High Edit Distance. Grows from 40.26% to 95.21%, reflecting surface mismatch.
5.2 Error Analysis
We categorize observed LLM failures into four major types, each of which presents recurring challenges as question complexity increases. For detailed error analysis on examples from each category, see Appendix E, Table 15.
Misinterpretation of Semantics LLMs frequently misinterpret task-specific terms such as “location type" or “position". Rather than grounding responses in document context, models often default to general world knowledge (e.g., answering “city" for a location type) or paraphrase the question, indicating weak comprehension of domain-specific semantics.
Implicit Information Gaps The model struggles to extract answers that are not explicitly stated but implied through document structure, such as signature blocks or formatting cues. For instance, it fails to associate a person with an organization unless the relationship is stated verbatim, ignoring visual or positional indicators that a human reader would readily infer.
Set Operation Failures To examine the impact of the number of set operations on model performance, we conducted an ablation experiment by varying one dimension within our three-dimensional parameter space while holding the others constant. Specifically, we increased the number of set operations while keeping the plurality and hop dimensions fixed. We observe that an increase in the number of set operations consistently resulted in a decline in performance metrics. Questions involving comparisons, intersections, or exclusions across entities consistently yield low F1 scores. The model fails to isolate shared or distinct roles when multiple organizations or individuals are involved, indicating limited capacity for multi-hop or set-based reasoning.
Long-Context Retrieval Errors Despite the presence of answers in the document, the model often returns “Not found", especially when relevant information appears in the middle of the context. This aligns with the “lost-in-the-middle" phenomenon (Liu et al., 2024b), where models disproportionately focus on the beginning or end of long texts, neglecting intermediate content.
Our error analysis reveals that while LLMs perform adequately on simple, single-step queries, they falter on multi-hop, set-operation, and implicit-knowledge tasks – particularly those requiring retrieval of details buried in the “middle" of the context. Addressing these shortcomings will likely involve a combination of targeted fine-tuning, improved context chunking or retrieval modules, and stronger in-context reasoning prompts.
6 Discussion & Conclusion
Our framework integrates structured annotations, knowledge graph construction, and templated QA extraction (Sections 3.2-3.5) to produce a dataset with finely controlled question complexity over real-world long document. This allows systematic evaluation across various dimensions with identifying nuanced model failures. Despite high scores on simple factual questions, even the strongest LLMs struggle with compositional logical inference, especially set-based comparisons and multi-hop paths, frequently failing to interpret implicit relationships and falling back on world knowledge instead of grounding answers in context (Section 5).
Applicability and Generalizability
As discussed in Sections 1 and 3.1, the ability to process long documents is essential for real-world QA tasks. Our benchmark illustrates this challenge using credit agreements, where information is densely structured, scattered, and cross-referenced. However, legal contracts, policy documents, patent filings, and other non-financial documents all exhibit similarly complex entity structures. Even in non-legal domains, such as medical records or scientific articles, multi-hop relationships and compositional multi-hop inference over entities are common, making our methodology broadly applicable.
Opportunities and Extensions
While our benchmark focuses on pre-trained LLMs, the structured design of KG-QAGen enables future adaptation, such as domain-specific fine-tuning. Expanding the dataset beyond U.S.-based financial documents could support geographically diverse evaluation.
Appendix A Annotation Guide
A.1 Labels
Our annotation framework captures key entities and relationships within credit agreements, focusing on seven labels that define organizational structure, roles, and locations.
Organization Name
The name of a company or financial institution mentioned in the agreement. Examples: "Goldman Sachs", "JP Morgan Chase".
Organization Role
The role that an organization holds in the agreement. Examples: "Lender", "Borrower", "Guarantor".
Organization Sub-Role
A more specific classification within an Organization Role that provides additional detail. Examples: "Lead Lender", "Administrative Agent".
Person Name
The name of an individual explicitly mentioned in the agreement. Examples: "John Doe", "Jane Smith".
Person Position
The title or role held by a person within an organization. Examples: "Vice President", "Loan Officer".
Location
A physical address associated with an organization. Examples: "35 Peachtree St, Atlanta, GA".
Location Type
A classification of a Location, providing additional information on its function. Example: "Headquarters", "Branch Office".
A.2 General rules
Unique Entity Annotation
Organization Names, Person Names, and Locations are annotated only once per document, even if they appear multiple times. Organization Roles, Organization Sub-Roles, Person Positions, and Location Types are annotated once for each unique entity they describe.
Example: In Figure 5(a), "BANK OF AMERICA, N.A." appears multiple times but is annotated only once. However, its roles, "Administrative Agent" and "Lender", are each annotated where they are explicitly stated. If another organization is a lender and/or an administrative agent, those role labels will be annotated as close as possible to the name label. Similarly for "Alexandra M. Knights" and her position "Authorized Signatory".
Ordered Search for Annotations
To maintain consistency, we follow a structured order when searching for annotations.
  • Signature Box: The signature box is the primary source for annotations, as it usually contains the most organizations (e.g., Parent Borrowers, Lead Lenders) and their roles. In most cases, organizations are explicitly labeled with their roles within this section, making it the most reliable source.
  • Headers: If an entity is missing in the signature box, we check the document headers. It is common for certain organizations’ roles to be absent from the signature box but present in section headings. In some cases, even company names are missing from the signature box and must be retrieved from headers.
  • Full Document: Locations are often not explicitly mentioned in the signature box or headers, requiring a broader document search. Occasionally, some organizations (e.g., Guarantors) are not mentioned in either the signature box or the headers, making a full-document search necessary.
Relation Annotations
  • Org Name $\rightarrow$ Org Role: Each organization name is annotated once per document, while its role is annotated as close to the name as possible. If multiple organizations share the same role (e.g., Borrower) and it is mentioned collectively in the signature box, the role is annotated once and linked to all relevant organizations. Otherwise, each occurrence of a role is annotated separately and linked to the corresponding organization.
  • Org Role $\rightarrow$ Org Sub-Role: If an organization has a specific sub-role (e.g., Parent Borrower), its role is annotated separately from others, and the sub-role is linked to it. The connection follows a hierarchical structure, linking the organization to its role, then the role to its sub-role.
  • Org Name $\rightarrow$ Person Name: Each person’s name is annotated once per document and linked to a single organization (see A.3 for special scenarios).
  • Person Name $\rightarrow$ Person Position: A person’s position is annotated as close to their name as possible. Individuals may hold multiple positions, which are all linked to the same person annotation. In some cases, position mentions may be missing (see Figure 5(d) and A.3).
  • Org Name $\rightarrow$ Location: Locations are annotated once per document and linked to the relevant organization.
  • Location $\rightarrow$ Location Type: Each location has exactly one type, which is annotated as close to the location name as possible. Location types are often missing, and in such cases, no annotation is applied (see Fig 5(d)).
Continuation Links
In cases where the text to be annotated is split across multiple lines or interrupted by symbols, preventing a single coherent annotation, we annotate each segment separately and link them sequentially: the first to the second, the second to the third, and so on. This issue is particularly common for Locations, where addresses or names are fragmented (see Figure 5(d)). Since our annotation structure does not include relations between labels of the same type, these links are easily identifiable. During post-processing, entities connected by continuation links are automatically merged with spaces, ensuring a single coherent label.
A.3 Main Edge Cases
Parent Borrower and Related Entities
Some agreements list multiple entities as borrowers and guarantors, often sharing highly similar names (e.g., Summit Hospitality JV, LP, Summit JV MR 2, LLC in Figures 5(b) and 5(c)), suggesting a structured division of what may functionally be a single entity. In such cases, one entity is designated as the Parent Borrower, while the others are categorized under Borrowers and Guarantors. The same individual typically signs on behalf of all these entities, but following our general rule, we annotate a person once and link them to the Parent Borrower. Additionally, since the Parent Borrower’s role is often labeled simply as Parent in the signature box, while other borrowers explicitly receive the Borrower role, we annotate the closest explicit mention of Borrower elsewhere in the document to ensure accurate labeling.
Missing Person Position
In some cases, individuals are mentioned within address sections rather than as signatories or officials, likely indicating their responsibility for a specific location. These individuals often lack an explicitly stated position. As shown in Figure 5(d), Jim Plocica appears in the address block without a corresponding title. When this occurs, we annotate the Person Name but leave the Person Position unannotated.
Additionally, this example highlights several annotation conventions:
  • The Organization Name (Bright Horizons Family Solutions LLC) is omitted here because it is annotated in the signature box (once per document). The links to the Location and Person Name originate from that primary annotation.
  • The Location spans multiple lines, requiring the use of Continuation Links (described in Section A.2) to maintain coherence.
  • No Location Type is explicitly mentioned, which is a common occurrence in such address sections.
Shared Roles and Prefixes
When multiple organizations share a role, they are often referred to collectively using prefixes like Joint or Co- (e.g., Joint Bookrunner or Co-Documentation Agent). However, our annotation framework requires capturing role information at the individual organization level. Since a single entity cannot hold a joint role independently, we do not annotate these prefixes. Instead, we annotate only the core role (Bookrunner, Documentation Agent, etc.) for each company separately, as illustrated in Figure 5(e)).
Additionally, in this example, the annotations are sourced from the header rather than the signature box because Merrill Lynch, Pierce, Fenner & Smith Incorporated is absent from the signature box, and for other companies, only their primary role (Lender) is mentioned there.
Smallest Document
One of the extracted credit agreements has only 2 organizations involved, and their roles are mentioned neither in the signature box nor in the header (Figure 5(f)). Moreover, the role of the lender is not mentioned anywhere in the document, while the role of the borrower is mentioned once (Figure 5(g)).
(a) General and most common annotation in the signature box.
(b) Multiple organizations with similar name being borrowers.
(c) Multiple organizations with similar name being guarantors in addition to being borrowers.
(d) Location without type, person without position and continuation links.
(e) Annotations in the header and joint roles.
(f) Only 2 organizations, not roles in the signature box or header.
(g) Only 2 organizations, only borrower’s role is mentioned.
Figure 5: Representative Annotation Scenarios from Label Studio.
Appendix B Knowledge Graph Construction Details
Ontology Layer
The ontology layer defines a class hierarchy that structures entities in the knowledge graph. We organize annotated entity types into three top-level classes — Person, Organization, and Location — based on their functional roles in financial agreements. These are connected to their respective sub-categories through explicit Sub-class relationships, capturing hierarchical distinctions. For example, Organization includes roles such as Lender, Agent, and L/C Issuer, while Person includes positions like Vice President and Treasurer. To further refine responsibilities, we introduce a second level of subclasses via Sub-Roles, such as dividing Agent into Administrative Agent and Documentation Agent. Only these role and position types appear as classes in the ontology layer; the actual named entities (e.g., “Jane Smith", “ABCDE Inc.") are treated as instances in the data layer. This abstraction supports generalization across documents and enables scalable, template-based QA pairs extraction.
Data Layer
Complementing the ontology layer, the data layer grounds the class structure in real-world instances and their relationships. Entities such as ABCDE Inc., John Doe, and 10 Peachtree instantiate ontology classes like Organization, Person, and Location via Instance Of links. These instances are further connected through Inter-Instance relations (e.g., hasEmployee, isLocationOf), capturing document-specific associations such as employment and location.
Ontology and Data Scope
While each credit agreement yields its own data layer containing document-specific instances and relationships, the ontology layer is shared across all documents. This unified ontology defines a consistent schema for entity types and roles, enabling systematic alignment across diverse agreements.
Question Construction with SPARQL
SPARQL queries enable automated, structured QA extraction from our knowledge graph. This approach dynamically adapts to diverse financial agreements by leveraging ontology. The query in Figure 6 retrieves individuals and their positions using the isInstanceOf relationship, while GROUP_CONCAT consolidates multiple positions per individual for comprehensive QA pairs extraction.
SELECT DISTINCT ?person (GROUP_CONCAT(?position; separator="|") as ?positions)
WHERE {
?person a <http://example.org/base/Person> ;
<http://example.org/isInstanceOf/> ?position .
FILTER(STRSTARTS(STR(?position), "http://example.org/person_position/"))
GROUP BY ?person
Figure 6: SPARQL Query Example for Structured QA Extraction.
Appendix C All templates
Table 5: Level 1 Question Templates.
Template Based On P H #SO
What is the position of [Person Name]? [if one] 0 1 0
In what organization does [Person Name] work? 0 1 0
Who is the representative of [Org Name]? [if one] 0 1 0
What is the role of [Org Name] in the agreement? [if one] 0 1 0
What company is the [Org Role (+ Sub-Role)] in the agreement? [if one] 0 1 0
What is the location of [Org Name]? [if one] 0 1 0
Which company is associated with [Location]? 0 1 0
What type of location is [Location] (e.g., Headquarters, Trade Operations, etc.)? [if one] 0 1 0
Table 6: Level 2 Question Templates.
Template Based On P H #SO
What are the positions of [Person Name]? What is the position of [Person Name]? [if one] 1 1 0
Who are the representatives of [Org Name]? Who is the representative of [Org Name]? [if one] 1 1 0
What are the roles of [Org Name] in the agreement? What is the role of [Org Name] in the agreement? [if one] 1 1 0
What companies are the [Org Role (+ Sub-Role)] in the agreement? What company is the [Org Role (+ Sub-Role)] in the agreement? [if one] 1 1 0
What are the locations of [Org Name]? What is the location of [Org Name]? [if one] 1 1 0
What types of location is [Location] (e.g., Headquarters, Trade Operations, etc.)? What type of location is [Location] (e.g., Headquarters, Trade Operations, etc.)? [if one] 1 1 0
What is the position held by both [Person Name 1] and [Person Name 2]? [if one] What is the position of [Person Name]? [if one] 0 1 1
What role do both [Org Name 1] and [Org Name 2] have in the agreement? [if one] What is the role of [Org Name] in the agreement? [if one] 0 1 1
What company is both the [Org Role (+ Sub-Role) 1] and [Org Role (+ Sub-Role) 2] in the agreement? [if one] What company is the [Org Role (+ Sub-Role)] in the agreement? [if one] 0 1 1
Who is the [Person Position] of [Org Name]? [if one] 0 2 0
What is the role in the agreement of the company where [Person Name] is employed? [if one] What is the role of [Org Name] in the agreement? [if one] 0 2 0
What is the role in the agreement of the company associated with [Location]? [if one] What is the role of [Org Name] in the agreement? [if one] 0 2 0
What is the [Location Type] office of [Org Name]? 0 2 0
Table 7: Level 3 Question Templates.
Template Based On P H #SO
What are the positions held by both [Person Name 1] and [Person Name 2]? What are the positions of [Person Name]? What is the position held by both [Person Name 1] and [Person Name 2]? [if one] 1 1 1
What roles do both [Org Name 1] and [Org Name 2] have in the agreement? What are the roles of [Org Name] in the agreement? What role do both [Org Name 1] and [Org Name 2] have in the agreement? [if one] 1 1 1
What companies are both the [Org Role (+ Sub-Role) 1] and [Org Role (+ Sub-Role) 2] in the agreement? What companies are the [Org Role (+ Sub-Role)] in the agreement? 1 1 1
What is the position held by [Person Name 1] but not by [Person Name 2]? [if one] What is the position held by both [Person Name 1] and [Person Name 2]? [if one] 0 1 2
What role does [Org Name 1] have in the agreement which is not the role of [Org Name 2]? [if one] What role do both [Org Name 1] and [Org Name 2] have in the agreement? [if one] 0 1 2
What company is the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] in the agreement? [if one] What company is both the [Org Role (+ Sub-Role) 1] and [Org Role (+ Sub-Role) 2] in the agreement? [if one] 0 1 2
Who are the [Person Position]s of [Org Name]? Who is the [Person Position] of [Org Name]? [if one] 1 2 0
What are the roles in the agreement of the company where [Person Name] is employed? What is the role in the agreement of the company where [Person Name] is employed? [if one] 1 2 0
What are the roles in the agreement of the company associated with [Location]? What is the role in the agreement of the company associated with [Location]? [if one] 1 2 0
Who is the [Person Position 1] and [Person Position 2] of [Org Name]? [if one] Who is the [Person Position] of [Org Name]? [if one] 0 2 1
Who is the [Person Position] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [if one, and the company should be uniquely identifiable] Who is the [Person Position] of [Org Name]? [if one] 0 3 0
Who is the [Person Position] of the company associated with [Location]? [if one] Who is the [Person Position] of [Org Name]? [if one] 0 3 0
Who is the [Person Position] of the company associated where [Person Name] is employed? [if one] Who is the [Person Position] of [Org Name]? [if one] 0 3 0
What is the address of [Location Type] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [if one, and the company should be uniquely identifiable] What is the [Location Type] office of [Org Name]? 0 3 0
What is the address of [Location Type] of the company associated with [Location (different)]? What is the [Location Type] office of [Org Name]? 0 3 0
What is the address of [Location Type] of the company where [Person Name] is employed? What is the [Location Type] office of [Org Name]? 0 3 0
Table 8: Level 4 Question Templates, part 1.
Template Based On P H #SO
What are the positions held by [Person Name 1] but not by [Person Name 2]? What are the positions held by both [Person Name 1] and [Person Name 2]? What is the position held by [Person Name 1] but not by [Person Name 2]? [if one] 1 1 2
What roles does [Org Name 1] have in the agreement which are not the roles of [Org Name 2]? What roles do both [Org Name 1] and [Org Name 2] have in the agreement? What role does [Org Name 1] have in the agreement which is not the role of [Org Name 2]? [if one] 1 1 2
What companies are the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] in the agreement? What companies are both the [Org Role (+ Sub-Role) 1] and [Org Role (+ Sub-Role) 2] in the agreement? What company is the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] in the agreement? [if one] 1 1 2
What is the position held by [Person Name 1] but not by [Person Name 2] or [Person Name 3]? [if one] What is the position held by [Person Name 1] but not by [Person Name 2]? [if one] 0 1 3
What is the position held by [Person Name 1] and [Person Name 2] but not by [Person Name 3]? [if one] What is the position held by [Person Name 1] but not by [Person Name 2]? [if one] 0 1 3
What role do [Org Name 1] and [Org Name 2] have in the agreement which is not the role of [Org Name 3]? [if one] What role does [Org Name 1] have in the agreement which is not the role of [Org Name 2]? [if one] 0 1 3
What role does [Org Name 1] have in the agreement which is not the role of [Org Name 2] or [Org Name 3]? [if one] What role does [Org Name 1] have in the agreement which is not the role of [Org Name 2]? [if one] 0 1 3
What company is the [Org Role (+ Sub-Role) 1] and [Org Role (+ Sub-Role) 2] but not the [Org Role (+ Sub-Role) 3] in the agreement? [if one] What company is the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] in the agreement? [if one] 0 1 3
What company is the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] [Org Role (+ Sub-Role) 3] in the agreement? [if one] What company is the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] in the agreement? [if one] 0 1 3
Table 9: Level 4 Question Templates, part 2.
Template Based On P H #SO
Who are the both [Person Position 1]s and [Person Position 2]s of [Org Name]? Who are the [Person Position]s of [Org Name]? Who is the [Person Position 1] and [Person Position 2] of [Org Name]? [if one] 1 2 1
Who are the [Person Position]s of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [the company should be uniquely identifiable] Who is the [Person Position] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [if one, and the company should be uniquely identifiable] 1 3 0
Who are the [Person Position]s of the company associated with [Location]? Who is the [Person Position] of the company associated with [Location]? [if one] 1 3 0
Who are the [Person Position]s of the company associated where [Person Name] is employed? Who is the [Person Position] of the company associated where [Person Name] is employed? [if one] 1 3 0
Who is the [Person Position] of the company which is both the [Org Role(-s) (+ Sub-Role(-s)) 1] and the [Org Role(-s) (+ Sub-Role(-s)) 2] in the agreement? [if one, and the company should be uniquely identifiable] Who is the [Person Position] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [if one, and the company should be uniquely identifiable] 0 3 1
Who is both the [Person Position 1] and [Person Position 2] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [if one, and the company should be uniquely identifiable] Who is the [Person Position] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [if one, and the company should be uniquely identifiable] 0 3 1
Who is both the [Person Position 1] and [Person Position 2] of the company associated with [Location]? [if one] Who is the [Person Position] of the company associated with [Location]? [if one] 0 3 1
Who is both the [Person Position 1] and [Person Position 2] of the company associated where [Person Name] is employed? [if one] Who is the [Person Position] of the company associated where [Person Name] is employed? [if one] 0 3 1
What is the address of the [Location Type] office of the company which is both the [Org Role(-s) (+ Sub-Role(-s)) 1] and the [Org Role(-s) (+ Sub-Role(-s)) 2] in the agreement? [if one, and the company should be uniquely identifiable] What is the [Location Type] office of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [if one, and the company should be uniquely identifiable] 0 3 1
Table 10: Level 5 Question Templates, part 1.
Template Based On P H #SO
What are the positions held by [Person Name 1] but not by [Person Name 2] or [Person Name 3]? What are the positions held by [Person Name 1] but not by [Person Name 2]? What is the position held by [Person Name 1] but not by [Person Name 2] or [Person Name 3]? [if one] 1 1 3
What are the positions held by [Person Name 1] and [Person Name 2] but not by [Person Name 3]? What are the positions held by [Person Name 1] but not by [Person Name 2]? What is the position held by [Person Name 1] and [Person Name 2] but not by [Person Name 3]? [if one] 1 1 3
What roles do [Org Name 1] and [Org Name 2] have in the agreement which are not the roles of [Org Name 3]? What roles does [Org Name 1] have in the agreement which are not the roles of [Org Name 2]? What role do [Org Name 1] and [Org Name 2] have in the agreement which is not the role of [Org Name 3]? [if one] 1 1 3
What roles does [Org Name 1] have in the agreement which are not the roles of [Org Name 2] or [Org Name 3]? What roles does [Org Name 1] have in the agreement which are not the roles of [Org Name 2]? What role does [Org Name 1] have in the agreement which is not the role of [Org Name 2] or [Org Name 3]? [if one] 1 1 3
What companies are the [Org Role (+ Sub-Role) 1] and [Org Role (+ Sub-Role) 2] but not the [Org Role (+ Sub-Role) 3] in the agreement? What companies are the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] in the agreement? What company is the [Org Role (+ Sub-Role) 1] and [Org Role (+ Sub-Role) 2] but not the [Org Role (+ Sub-Role) 3] in the agreement? [if one] 1 1 3
What companies are the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] [Org Role (+ Sub-Role) 3] in the agreement? [if one] What companies are the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] in the agreement? What company is the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] [Org Role (+ Sub-Role) 3] in the agreement? [if one] 1 1 3
Table 11: Level 5 Question Templates, part 2.
Template Based On P H #SO
Who are the [Person Position]s of the company which is both the [Org Role(-s) (+ Sub-Role(-s)) 1] and the [Org Role(-s) (+ Sub-Role(-s)) 2] in the agreement? [the company should be uniquely identifiable] Who are the [Person Position]s of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [the company should be uniquely identifiable] Who is the [Person Position] of the company which is both the [Org Role(-s) (+ Sub-Role(-s)) 1] and the [Org Role(-s) (+ Sub-Role(-s)) 2] in the agreement? [if one, and the company should be uniquely identifiable] 1 3 1
Who are both the [Person Position 1] and [Person Position 2] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [the company should be uniquely identifiable] Who are the [Person Position]s of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [the company should be uniquely identifiable] Who is both the [Person Position 1] and [Person Position 2] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [if one, and the company should be uniquely identifiable] 1 3 1
Who are both the [Person Position 1] and [Person Position 2] of the company associated with [Location]? Who are the [Person Position]s of the company associated with [Location]? Who is both the [Person Position 1] and [Person Position 2] of the company associated with [Location]? [if one] 1 3 1
Who are both the [Person Position 1] and [Person Position 2] of the company associated where [Person Name] is employed? Who are the [Person Position]s of the company associated where [Person Name] is employed? Who is both the [Person Position 1] and [Person Position 2] of the company associated where [Person Name] is employed? [if one] 1 3 1
Who is the [Person Position 1] but not [Person Position 2] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [if one, and the company should be uniquely identifiable] Who is both the [Person Position 1] and [Person Position 2] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [if one, and the company should be uniquely identifiable] 0 3 2
Who is the [Person Position 1] but not [Person Position 2] of the company associated with [Location]? [if one] Who is both the [Person Position 1] and [Person Position 2] of the company associated with [Location]? [if one] 0 3 2
Who is the [Person Position 1] but not [Person Position 2] of the company associated where [Person Name] is employed? [if one] Who is both the [Person Position 1] and [Person Position 2] of the company associated where [Person Name] is employed? [if one] 0 3 2
What is the [Location Type] office of the company which is both the [Org Role(-s) (+ Sub-Role(-s)) 1] but not the [Org Role(-s) (+ Sub-Role(-s)) 2] in the agreement? [if one, and the company should be uniquely identifiable] What is the [Location Type] office of the company which is both the [Org Role(-s) (+ Sub-Role(-s)) 1] and the [Org Role(-s) (+ Sub-Role(-s)) 2] in the agreement? [if one, and the company should be uniquely identifiable] 0 3 2
Appendix D Evaluation Details
D.1 Prompting Strategy
Batch QA Prompt [SYSTEM INPUT]
You are a financial expert, and your task is to answer the questions given to you in batches about the provided credit agreement. If you believe the answer is not present in the agreement, say ’Not found’.
[EXPECTED OUTPUT]
Respond ONLY with valid JSON, nothing else. See the example below.
The given document:
Tesla, Inc. is an American electric vehicle and clean energy company founded in 2003 by Martin Eberhard and Marc Tarpenning. Elon Musk became the largest investor and later CEO.
The given questions:
Q1: Who founded Tesla?
Q2: What year was Tesla founded?
The expected output:
    "answers": [               {"question_index": 1,               "answer": "Martin Eberhard,                                 Marc Tarpenning"},               {"question_index": 2,               "answer": "2003"}        ] }
[USER INPUT]
{Document Context}
[QUESTIONS]
Q1: {question}
Q2: {question}
(a) Example of a batch prompt combining up to 50 questions with a single document chunk. The LLM is instructed to return only JSON-formatted answers, each linked to a question index.
Merging Prompt [SYSTEM INPUT]
You are a financial expert, and your task is to combine or merge the provided partial answers, coming from different chunks of a credit agreement, into a single final answer for each of the questions given to you. If you believe the answer is not present in the agreement, say ’Not found’.
[EXPECTED OUTPUT]
The given document:
Tesla, Inc. is an American electric vehicle and clean energy company founded in 2003 by Martin Eberhard and Marc Tarpenning. Elon Musk became the largest investor and later CEO.
The given questions:
Q1: Who founded Tesla?
Q2: What year was Tesla founded?
The expected output:
    "answers": [               {"question_index": 1,               "answer": "Martin Eberhard,                                 Marc Tarpenning"},               {"question_index": 2,               "answer": "2003"}        ] }
[USER INPUT]
Chunk 1 partial answer JSON: {ans 1}
Chunk 2 partial answer JSON: {ans 2}
[QUESTIONS]
Q1: {question}
Q2: {question}
(b) Example of a merging prompt that consolidates partial answers from multiple document chunks into one final JSON-formatted answer per question.
Figure 7: Two-stage prompting pipeline. Long documents are chunked if needed and paired with question batches ($\leqslant\!50$). A structured one-shot prompt is used per chunk. For multi-chunk documents, a merging prompt consolidates partial answers.
We implement a two-stage prompting pipeline for QA inference over long documents. In the first stage, the document is either processed in full or split into chunks (depending on the context limit of the LLM), and each chunk is paired with a batch of up to 50 questions. We then apply a one-shot, instruction-based prompt (Figure 7(a)) that instructs the LLM to act as a financial expert and return an array of answers – either grounded in the document or explicitly stating “Not found" – in valid JSON format.
The expected format is enforced both through the prompt instructions and through structured parsing logic. The parser first attempts to extract a JSON object from the model’s response, supporting both raw JSON and code-fenced output. If parsing fails, the system falls back to extracting answers from list-formatted outputs or reattempts the query, up to a maximum number of retries. All models are queried with temperature=0.0 for determinism.
In the second stage, for questions where answers may be distributed across multiple chunks, a merging prompt is issued (Figure 7(b)). This prompt consolidates partial answers from earlier responses and asks the model to output a single final answer per question, again using the same strict JSON format. This strategy allows the model to incorporate information across multiple views of the document.
Regarding resources, as mentioned in 4.1, for open-source models we are using Together.AI for some of the models. For the other open-source models, we used 1 H200 GPU.
D.2 Evaluation Metrics
We employ four complementary metrics to evaluate model responses.
  • Word-Level F1 Score  Token-level precision and recall against the gold answer, ensures strict factual correctness in extractive question answering.
  • Normalized Levenshtein Distance  Character-level string distance, indicating surface-level divergence (described by Yujian and Bo (2007)).
  • Cosine Similarity  Evaluates the semantic alignment, providing robustness to minor changes.
  • LLM-as-a-Judge  Gemini-2.0-Pro is prompted to score each prediction from 1 to 5 based on its semantic correctness, given the question and reference answer.
We exclude ROUGE, BLEU, and similar metrics because they are known to perform poorly on short factual answers, tend to over-penalize surface-level variations and are not well-aligned with human judgment in our setting (Blagec et al., 2022; Chaganty et al., 2018).
Table 12: Comparison of model performance across difficulty levels, including F1 Score, Normalized Edit Distance, Cosine Similarity, and LLM-as-a-Judge. denotes the models evaluated on a smaller subset due to resources constraints.
Easy Medium Hard
Model F1 Edit Cosine LLM-as- F1 Edit Cosine LLM-as- F1 Edit Cosine LLM-as-
Score $\uparrow$ Dist. $\downarrow$ Sim. $\uparrow$ a-Judge $\uparrow$ Score $\uparrow$ Dist. $\downarrow$ Sim. $\uparrow$ a-Judge $\uparrow$ Score $\uparrow$ Dist. $\downarrow$ Sim. $\uparrow$ a-Judge $\uparrow$
Proprietary LLMs
Gemini-2.0-Flash 0.6748 0.2679 0.6650 4.0000 0.3294 0.6280 0.3121 2.4820 0.1172 0.7238 0.0608 2.2838
Gemini-2.0-Flash-Lite 0.5908 0.3763 0.5829 3.4172 0.3883 0.5985 0.3715 2.6497 0.1943 0.7956 0.1255 1.5135
Gemini-1.5-Pro 0.5410 0.3903 0.5331 3.2448 0.4577 0.5340 0.4415 2.8436 0.1791 0.7786 0.1780 1.6986
Gemini-1.5-Flash 0.5103 0.4285 0.5032 3.1207 0.3892 0.6101 0.3657 2.5130 0.1400 0.8389 0.1255 1.5411
GPT-4o 0.5272 0.4177 0.5213 3.2034 0.3783 0.6098 0.3690 2.5451 0.1178 0.7738 0.1144 1.4795
GPT-4o-Mini 0.4723 0.4614 0.4654 2.9690 0.2732 0.6756 0.2740 2.0681 0.0274 0.8294 0.0336 1.1370
Open-source LLMs
DeepSeek-V3 0.6218 0.3722 0.6077 3.5668 0.4796 0.4912 0.4678 3.0210 0.0856 0.7975 0.0737 1.4795
Llama-4-Maverick-Instruct 0.5552 0.3585 0.5494 3.4482 0.3848 0.5606 0.3669 2.6976 0.2031 0.7231 0.1172 3.1351
Llama-4-Scout-Instruct 0.4993 0.4114 0.4920 3.2138 0.2215 0.6765 0.2099 1.8772 0.0472 0.8225 0.0303 1.5270
Llama-3.1-Instruct-Turbo 0.5472 0.3901 0.5389 3.4793 0.4204 0.5489 0.4018 2.6377 0.3651 0.7197 0.2849 2.7297
Llama-3.1-Instruct 0.4955 0.4707 0.4847 3.1373 0.4017 0.5481 0.3917 2.6722 0.1639 0.7667 0.1106 2.3425
Llama-3.1-Instruct 0.3891 0.5617 0.3793 2.6996 0.2962 0.6255 0.2856 2.2332 0.1169 0.7980 0.0915 1.4315
Qwen2-Instruct 0.3886 0.6369 0.3710 3.3884 0.3100 0.7031 0.2986 2.9078 0.0762 0.8200 0.0690 1.3288
D.3 Comparison with Oracle and RAG
We compare three retrieval settings for answering questions from credit agreements: (1) Oracle, which uses the pieces of text containing the answer; (2) RAG (Retrieval-Augmented Generation), which selects top passages using a retriever; and (3) Full, where the entire agreement is provided to the model. In table 13, we observe that RAG performs drastically worse than the other two settings across all difficulty levels, which highlights the incompatibility of standard retrieval techniques with dense, cross-referenced financial documents, often requiring the aggregation of logically related but spatially distant elements.
Interestingly, the comparison between Oracle and Full Document settings reveals a non-trivial pattern. For easy look-up questions, the Oracle setting performs better, which aligns with expectations since such questions benefit from isolating the minimal, most relevant snippet. However, for medium and hard questions, the Full Document consistently outperforms the Oracle setting. This suggests that isolating the “correct" snippet removes necessary contextual cues that models need for logical, multi-hop inference. These findings reinforce the value of credit agreements as a testbed for long-context LLMs, and demonstrate how our framework’s question complexity help uncover the precise scenarios where models struggle, even when given the exact ground-truth passage (see Section 5 and E for the detailed analysis).
Table 13: Performance comparison across three retrieval settings – Oracle (gold snippet only), RAG (retriever-selected passages), and Full (entire credit agreement) – for each question complexity group.
Easy Medium Hard
Setting F1 Edit Cosine LLM-as- F1 Edit Cosine LLM-as- F1 Edit Cosine LLM-as-
Score $\uparrow$ Dist. $\downarrow$ Sim. $\uparrow$ a-Judge $\uparrow$ Score $\uparrow$ Dist. $\downarrow$ Sim. $\uparrow$ a-Judge $\uparrow$ Score $\uparrow$ Dist. $\downarrow$ Sim. $\uparrow$ a-Judge $\uparrow$
Full 0.8542 0.1252 0.8507 4.4893 0.6936 0.3417 0.6932 3.8983 0.1172 0.7238 0.0608 2.2838
Oracle 0.9113 0.0673 0.9076 4.6950 0.6335 0.4075 0.6360 4.0508 0.1128 0.8067 0.0736 1.9189
RAG 0.1481 0.7341 0.1473 1.6099 0.1277 0.7543 0.1251 1.5254 0.0191 0.8450 0.0107 1.2027
D.4 Human Evaluation Metrics
We conducted a human evaluation of the LLM-generated responses using three independent evaluators. Two of these evaluators were annotators familiar with credit agreements, while the third evaluator had no prior exposure to the documents. Each evaluator was requested to rate the LLM responses on a scale from 1 to 5, according to the criteria below. A total of 100 questions were sampled from multiple templates to ensure maximal diversity in the human evaluation task.
  • 1   response is irrelevant or does not match the actual answer;
  • 2   response is somewhat relevant but does not match the actual answer;
  • 3   response shows slight similarity but lacks accuracy;
  • 4   response is largely relevant and accurate;
  • 5   response is a perfect match to the actual answer.
We assess the relationship between human evaluation scores and other metrics using both Pearson and Kendall’s Tau correlation coefficients (see Figure 8 and Table 14). Pearson correlation captures linear relationships, while Kendall’s Tau assesses rank-based similarity, making them complementary for evaluating metric alignment. Across both measures, F1 Score and LLM-as-a-Judge consistently show the highest positive correlations with average human ratings—indicating not only strong agreement in scale but also in ranking. This indicates that both F1 Score and LLM-as-a-Judge closely align with human evaluations, supporting their use as strong performance metrics.
Table 14: Kendall’s Tau and Pearson Correlation Tables.
Kendall’s Tau Correlation Matrix
F1 Score Cosine Similarity Edit Distance Sim LLM-as-a-Judge Avg LLM-Human Rating
F1 Score 1.000 0.902 -0.745 0.854 0.870
Cosine Similarity 0.902 1.000 -0.746 0.894 0.858
Edit Distance Sim -0.745 -0.746 1.000 -0.734 -0.733
LLM-as-a-Judge 0.854 0.894 -0.734 1.000 0.867
Avg LLM-Human Rating 0.870 0.858 -0.733 0.867 1.000
Pearson Correlation Matrix
F1 Score Cosine Similarity Edit Distance Sim LLM-as-a-Judge Avg LLM-Human Rating
F1 Score 1.000 0.933 -0.894 0.922 0.941
Cosine Similarity 0.933 1.000 -0.910 0.935 0.922
Edit Distance Sim -0.894 -0.910 1.000 -0.844 -0.862
LLM-as-a-Judge 0.922 0.935 -0.844 1.000 0.947
Avg LLM-Human Rating 0.941 0.922 -0.862 0.947 1.000
(a) Pearson correlation between various metrics. F1 Score and Avg LLM-Human Rating (r = 0.941), and LLM-as-a-Judge and Avg LLM-Human Rating (r = 0.947) show the highest correlations in the matrix.
(b) Kendall-Tau correlation between various metrics. The highest correlations are observed between F1 Score and Avg LLM-Human Rating ($\tau$ = 0.870), and between LLM-as-a-Judge and Avg LLM-Human Rating ($\tau$ = 0.867).
Figure 8: Kendall’s Tau and Pearson Correlation Matrices.
Appendix E Error Analysis
Below are detailed examples of errors from each of the aforementioned error categories.
Table 15: Error analysis with examples.
Question What type of location is Dubai, UAE?
Expected Answer Zonal Office
LLM Answer City
Error Explanation Misinterpretation of Semantics. The LLM misinterprets the semantic intent of the question. While City is factually accurate, the question requires identifying the Location Type of "Dubai, UAE" in the context of the credit agreement. The expected answer, zonal office, reflects a more precise understanding of this context.
Question In what organizations does Hamish Sandhu work?
Expected Answer DBG Holdings Subsidiary Inc, Differential Brands Group Inc
LLM Answer RG Parent LLC
Error Explanation Implicit Information Gaps. The LLM fails to infer implicit information due to structural complexity in the document. In cases involving multiple signatories or nested signature panels, human readers can intuitively associate individuals with their respective organizations based on layout and context. However, this relationship is not explicitly stated, and the LLM, lacking this inferential reasoning, incorrectly attributes Hamish Sandhu to the wrong organization.
Question What is the position held by Alexander Wu but not by Chwan Ming Ho or Edward Chen?
Expected Answer SVP
LLM Answer SVP and General Manager
Error Explanation Set Operation Failures. The LLM struggles with this query due to the complexity introduced by multiple set operations, in this case, requiring a comparison of roles across three individuals. Alexander Wu is the SVP and General Manager, while Chwan-Ming Ho and Edward Chen are both VPs and General Managers (see figure  9). While "SVP" is correct, the inclusion of "General Manager" indicates a failure to correctly subtract shared titles.
Question What is the role of Celanese Chemicals, Inc. in the agreement?
Expected Answer Subsidiary
LLM Answer Not found
Error Explanation Long-Context Retrieval Limitations. The model fails to retrieve the correct role because the organizational label "subsidiary" is not explicitly stated in a natural sentence. Instead, Celanese Chemicals, Inc. appears beneath a heading or clause indicating subsidiaries, often following the signature panel. This implicit semantic link requires structural and hierarchical reasoning that current LLMs struggle with—particularly when role information is implied by layout rather than expressed textually.
(a) Multiple entities and roles are listed in close proximity, making it difficult for the model to associate roles accurately.
(b) Alexander Wu
(c) Chwan Ming Ho
(d) Edward Chen
Figure 9: (a) In cases involving multiple signatories, the LLM cannot infer implicit information like Person Positions. (b), (c) and (d) illustrate the LLM’s inability to deal with multiple set operations.
Figure 10: Trends in Error Metrics Across Template Levels. The plots illustrate a clear upward trend across all error metrics as template complexity increases. This includes the percentage of LLM responses marked as "Not Found", as well as those exhibiting low F1 scores, low cosine similarity, and high edit distance – indicating a decline in accuracy with higher-level templates.
(a) Error metrics across set operations.
(b) Error metrics across number of hops.
Figure 11: Comparison of error metrics for Deepseek-V3. (a) Trends by Set Operations. (b) Trends by Number of Hops. With increase in number of hops or set operations, the percentage distribution of each error type increases.
Appendix F Dataset Release and Leaderboard Setup
Table 16: Dataset statistics of the constructed KG-QAGen-D dataset.
Stats Dev Test Total
# Documents 40 130 170
[…]
# Question per doc (avg) 14.75 23.49 21.44
# Question per doc (max) 83 428 428
[…]
# Medium Questions (count) 2,680 10,203 12,883
# Hard Questions (count) 239 467 706
[…]
To facilitate reproducibility and future research, we are publicly releasing (under CC-BY NC ND 4.0 license) QA pairs for 40 documents out of the total 170 as the development set, allowing researchers to validate model performance under controlled conditions. The remaining QA pairs serve as the test set, which is not publicly available to prevent data contamination (Sainz et al., 2023). Instead, we (1) are making the test set questions public, without the ground-truth answers, and (2) will host an online evaluation platform on HuggingFace [^6] https://huggingface.co/, enabling researchers to submit model predictions for the test set and to benchmark performance through a leaderboard. This setup follows standard practices in recent LLM evaluation benchmarks (Lu et al., 2023; Yue et al., 2024; Zhao et al., 2024a), ensuring that models are assessed under zero-shot conditions without access to test set ground-truth answers. The distribution of the test and dev samples is provided in Table 16.
Appendix G Ethics Statement
We generally avoid mentioning specific agreements and their details. Instead, we use placeholders such as ABCDE Inc. for organization names, Richard Roe for person names, and 10 Peachtree... for locations to illustrate the structure of agreement entities and their relations. However, in our annotation guide (Appendix A) and error analysis (Appendix E), we include screenshots of real document annotations and mention specific examples to highlight the complexity of these agreements. Moreover, we store the raw HTML files of the agreements in our GitHub repository, enabling other researchers to reproduce our benchmark or conduct further experiments. As these documents are publicly available through the SEC’s EDGAR system, their inclusion for both illustrative and experimental purposes raises no ethical concerns.
Appendix H Acknowledgments
We would like to thank Archishman V B, Jorge Navarro, Krish Shah, Kritika Bansal, Meher Bhardwaj, Nilesh G, Rushi Glasswala, Soumyajit Basu, Tvisha Shah and Xu Huang for contributing to the initial round of annotations and early-stage coding experiments.
Appendix I Author Contributions
NT led the data acquisition by extracting credit agreements from SEC EDGAR, developed the annotation guidelines specifying the entity and relation schemas, conceptualized the RDF-based knowledge graph framework, devised multi-dimensional QA templates, and benchmarked long-context LLMs in both RAG and oracle settings. VK extensively annotated and cleaned documents, identified edge cases and was involved in developing the KG construction and question-answer generation modules. She performed error analysis, led human evaluation, and constructed visualizations. HS contributed to the document annotation process, developed the pipelines for KG construction and SPARQL QA generation and assisted in LLM response analysis. AR implemented the initial end-to-end benchmarking pipeline and supported the human evaluation. He experimented with various prompting techniques and evaluation metrics and created one-shot prompts used in experiments. HSA contributed to all annotation stages and SPARQL-based question generation. VS, AL, and RL contributed to annotation, data cleaning, and manuscript refinement. AS supervised all iterations of the annotation process, guided edge-case incorporation, shaped research direction for document-level question answering, advised on ablation design, and contributed to error analysis and writing. SC provided overall project supervision and guidance.
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The last paragraph of introduction explicitly describes the contributions. The abstract conveys the same idea slightly differently but without overclaiming or contradicting.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: See paragraph “Opportunities and Extensions" in the Discussion & Conclusion section.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: No theoretical layouts, proofs, etc. are contributed.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We open-source the benchmarking codes and a part of the dataset, and mention it in the introduction, in the Dataset Release and Leaderboard Setup appendix. In addition, README file in our repositories provide explicit instructions to running the codes and getting the results.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We open-source only a part of the data, but we provide links to the data and the benchmarking codes in the header with instructions in the repositories. In additional, we provide reasoning to open-sourcing only a part of the data (see Appendix F Dataset Release and Leaderboard Setup).
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: See Appendix D.1 Prompting Strategy for the technical details of implementation, and Appendix F Dataset Release and Leaderboard Setup for the test-dev split statistics.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [N/A]
  34. Justification: The paper focuses on a evaluation of QA pairs, and treats it as deterministic rather than probabilistic, so error estimation is not discussed. Having said that, we conduct human evaluation of LLMs’ responses (on a sample) and calculate correlations with the other metrics to see which ones align with the human preferences better.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: See the last paragraph of appendix D.1 Prompting Strategy.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We believe our work fully conforms with the NeurIPS Code of Ethics. We also provide Ethics Statement (Appendix G) for clarification of using credit agreements.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: Our work does not introduce new model capabilities or deployment scenarios, so societal impacts are not directly applicable to the scope of this study.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We believe our benchmark does not have high risk of misuse since the source data is publicly available documents with no images, personal information, etc. We mention that we believe there are no ethics concerns in Appendix G Ethics Statement.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Everyone who contributed to the project till its final stage became a co-author. The other contributers are acknowledge in Appendix H Acknowledgments.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: Both repositories (for data and codes) have README files with instructions to replicate the benchmarking pipeline.
  64. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  65. 14.
    Crowdsourcing and research with human subjects
  66. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  67. Answer: [N/A]
  68. Justification: No crowdsourcing or research with human subjects was conducted.
  69. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  70. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  71. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  72. Answer: [N/A]
  73. Justification: No crowdsourcing or research with human subjects was conducted.
  74. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  75. 16.
    Declaration of LLM usage
  76. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  77. Answer: [N/A]
  78. Justification: LLMs were used only for refining the writing, not for a original contribution.
  79. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v2

@@
-KG-QAGen: A Knowledge-Graph-Based Framework for Systematic Question Generation and Long-Context LLM Evaluation
+KG-MuLQA: A Framework for KG-based Multi-Level QA Extraction and Long-Context LLM Evaluation
 Abstract
-The increasing context length of modern language models has created a need for evaluating their ability to retrieve and process information across extensive documents.
-While existing benchmarks test long-context capabilities, they often lack a structured way to systematically vary question complexity.
-We introduce KG-QAGen (Knowledge-Graph-based Question-Answer Generation), a framework that (1) extracts QA pairs at multiple complexity levels (2) by leveraging structured representations of financial agreements (3) along three key dimensions – multi-hop retrieval, set operations, and answer plurality – enabling fine-grained assessment of model performance across controlled difficulty levels.
-Using this framework, we construct a dataset of 20,139 QA pairs (the largest number among the long-context benchmarks) and open-source a part of it.
-We evaluate 13 proprietary and open-source LLMs and observe that even the best-performing models are struggling with set-based comparisons and multi-hop logical inference.
+We introduce KG-MuLQA (Knowledge-Graph-based Multi-Level Question-Answer Extraction): a framework that (1) extracts QA pairs at multiple complexity levels (2) along three key dimensions – multi-hop retrieval, set operations, and answer plurality, (3) by leveraging knowledge-graph-based document representations.
+This approach enables fine-grained assessment of model performance across controlled difficulty levels.
+Using this framework, we construct a dataset of 20,139 QA pairs based on financial credit agreements and evaluate 16 proprietary and open-weight Large Language Models, observing that even the best-performing models struggle with set-based comparisons and multi-hop reasoning over long contexts.
 Our analysis reveals systematic failure modes tied to semantic misinterpretation and inability to handle implicit relations.
+KG-MuLQA: A Framework for KG-based Multi-Level QA Extraction
+and Long-Context LLM Evaluation
+Nikita Tatarinov,🖂 Vidhyakshaya Kannan,* Haricharana Srinivasa,* Arnav Raj, Harpreet Singh Anand, Varun Singh, Aditya Luthra, Ravij Lade, Agam Shah,🖂 Sudheer Chava Georgia Institute of Technology 🖂 Corresponding Authors: {ntatarinov3, ashah482}@gatech.edu * Indicates equal contribution
 1 Introduction
-The increasing context length of recent language models has led to growing interest in evaluating their capabilities.
-Researchers have studied challenges like the Lost in the Middle problem (Liu et al., 2024b), examined the limiting factors of scaling models to long context (such as the constraints imposed by RoPE’s base value (Men et al., 2024)), and developed techniques to extend context length efficiently (Hooper et al., 2024; Ding et al., 2024).
-Others have focused on improving inference efficiency (Jiang et al., 2024b; Tang et al., 2024) or enhancing long-context utilization and mitigating information loss (Zhang et al., 2024d; Lin et al., 2024; Zhang et al., 2024c).
-Numerous benchmarks have been developed to evaluate long-context LLMs, covering a variety of tasks and capabilities.
-Some benchmarks support multi-lingual evaluation (Agrawal et al., 2024; Bai et al., 2024), include tasks such as segment sorting (Wang et al., 2024a; Dong et al., 2024), numerical reasoning (Zhao et al., 2024c; Reddy et al., 2024), timeline reordering (Zhang et al., 2024a).
-However, certain aspects of understanding long-context remain underexplored in existing benchmarks.
-Figure 1: Overview of KG-QAGen.
+Figure 1: Overview of KG-MuLQA.
 Credit agreements are annotated to identify entities and their relationships, forming a knowledge graph representation.
 This graph is then used to systematically extract multi-level QA pairs, which serve as the basis for benchmarking long-context LLMs.
-Real-World Documents
-Understanding large, real-world documents presents a unique challenge due to complex entity relationships, co-reference variations, and cross-referenced information scattered across sections (Zaporojets et al., 2022).
-This complexity makes credit agreements a precise fit for long-context evaluation, as they contain networks of entities referenced in varied ways across clauses (Section 3.1) and introduce numerous challenging edge cases (Appendix A).
-Question Complexity Levels
-Long-context benchmarks can offer deeper insight by distinguishing question difficulty levels, enabling analysis of how performance degrades with increasing logical inference depth – whether due to retrieval limits, context length, or multi-step inference.
-While some benchmarks contain tasks of varying difficulty (Gupta et al., 2024; Zhao et al., 2024c; Zhu et al., 2024), none explicitly categorize question complexity levels.
-Moreover, our benchmark not only defines these levels but also provides a systematic approach to creating granular questions, leveraging our annotations to deterministically construct QA pairs with clearly defined complexity directions.
-We introduce KG-QAGen (Knowledge-Graph-based Question-Answer Generation), a framework for structured extraction of QA pairs from manually annotated documents via knowledge-graph representations.
-Our approach structures questions along three dimensions: the number of hops, the use of set operations, and whether the answer is singular or plural.
-Using a knowledge-graph-inspired representation, we create diverse questions and extract answers efficiently with SPARQL queries.
-This structured approach ensures scalability, flexibility, and controlled dataset creation, enabling a more precise evaluation of long-context LLMs.
-We apply KG-QAGen to 170 credit agreements, resulting in KG-QAGen-D, a dataset of 20,139 QA pairs categorized by three complexity levels.
-We use it to evaluate 13 long-context LLMs and find that even on simple factual queries, some models fail to extract basic information; performance further drops with multi-hop paths and compositional questions involving set operations (Section 4.3).
-Our analysis shows that set-based comparisons are especially challenging, with high rates of “Not Found" responses and semantic misinterpretation (Section 5.2).
-Our contributions are as follows: (1) We present a framework grounded in knowledge graphs, enabling systematic extraction of QA pairs at varying complexity levels.
-(2) We open-source a portion of the KG-QAGen-D to encourage future work (see Appendix F for details).
-(3) We release our inference pipeline and benchmarking codes to facilitate reproducible evaluation of long-context models.
-2 Related Work
-Table 1: Comparison of KG-QAGen-D with existing long-context LLM benchmarks.
-Our benchmark stands out by evaluating question complexity on a granular level, incorporating real-world SEC credit agreements, and supporting multi-hop and context-aware retrieval.
-Real-world Multi-hop Context-aware Ques. complexity
-Dataset documents retrieval retrieval levels Document Type # Samples
-mLongRR (Agrawal et al., 2024) partially BBC News unspecified
-L-Eval (An et al., 2024) partially Earnings Calls + Misc 2,008
+Table 1: Comparison of KG-MuLQA-D with existing long-context LLM benchmarks.
+Multi-hop Reasoning marks whether the dataset requires reasoning across multiple evidence pieces; KG-MuLQA-D explicitly defines and scales multi-hop paths through its knowledge-graph structure, enabling controlled evaluation of relational inference depth.
+Systematic Q-Extraction denotes whether questions are generated deterministically rather than manually or by LLMs; KG-MuLQA-D automates QA generation from annotated graphs, preserving real document semantics without synthetic injection.
+Question Complexity shows whether the benchmark categorizes questions by difficulty; KG-MuLQA-D introduces a principled, multi-dimensional notion of complexity (hops + set operations + plurality), allowing granular performance analysis across reasoning levels.
+Long Multi-hop Systematic Question
+Dataset context reasoning Q-extraction complexity Document Type # Samples
+L-Eval (An et al., 2024) Earnings Calls + Misc 2,008
 CLongEval (Qiu et al., 2024) partially THUnews + Misc 7,267
+Marathon (Zhang et al., 2024a) partially Prior Datasets 1,530
+MMLongBench-DOC (Ma et al., 2024) partially arXiv + Misc 1,082
+DocFinQA (Reddy et al., 2024) partially SEC 10-K & 10-Q 7,437
 NOCHA (Karpinska et al., 2024) Fictional Books 2,002
-Marathon (Zhang et al., 2024a) Prior Datasets 1,530
+LooGLE (Li et al., 2024) arXiv + Misc 6,448
+BAMBOO (Dong et al., 2024) ACL Anthology + Misc 1,502
+LongBench (Bai et al., 2024) Previous Datasets 4,750
+Loong (Wang et al., 2024) SEC + Misc 1,600
+FinDVer (Zhao et al., 2024b) SEC 10-K & 10-Q 2,100
+M4LE (Kwan et al., 2024) partially Prior Datasets unspecified
+Michelangelo (Vodrahalli et al., 2024) partially partially synthetic unspecified
 […]
-M4LE (Kwan et al., 2024) partially News + Misc unspecified
-LooGLE (Li et al., 2024) partially arXiv + Misc 6,448
-BAMBOO (Dong et al., 2024) partially ACL Anthology + Misc 1,502
-LongBench (Bai et al., 2024) Previous Datasets 4,750
-MMLongBench-DOC (Ma et al., 2024) partially arXiv + Misc 1,082
-Loong (Wang et al., 2024b) SEC + Misc 1,600
-FinDVer (Zhao et al., 2024b) SEC 10-K & 10-Q 2,100
-DocFinQA (Reddy et al., 2024) SEC 10-K & 10-Q 7,437
+mLongRR (Agrawal et al., 2024) needles BBC News unspecified
 (Gupta et al., 2024) partially Financial News 560
+DocMath-Eval (Zhao et al., 2024c) partially partially SEC 10-K & 10-Q 4,000
 FanOutQA (Zhu et al., 2024) partially Wikipedia 7,305
-DocMath-Eval (Zhao et al., 2024c) partially SEC 10-K & 10-Q 4,000
-KG-QAGen-D (ours) SEC Credit Agreements 20,139
-Structured retrieval from long-context documents remains a key challenge, especially when handling complex entities and their relationships (Lin et al., 2025; Guo et al., 2024; Saad-Falcon et al., 2023).
-Multi-hop retrieval and context-aware models are particularly relevant for real-world documents.
-Multi-Hop Retrieval
-Multi-hop retrieval involves retrieving multiple interconnected pieces of evidence across different sections of a document or multiple documents to answer a question.
-This capability is crucial for handling complex queries requiring multi-step retrieval, such as tracing relationships between entities in legal documents or extracting financial conditions scattered across an agreement (Wang et al., 2025; Sarmah et al., 2024).
-Some existing benchmarks (An et al., 2024; Wang et al., 2024a) focus on long-context understanding without incorporating explicit multi-hop retrieval challenges.
-Context-Aware Retrieval
-Context-aware retrieval refers to tasks where models retrieve relevant information to answer questions based on document content rather than artificial information placement.
-Some existing benchmarks focus on the Needle in a Haystack problem (Agrawal et al., 2024), where small pieces of information are inserted into long texts and then retrieved.
-This type of retrieval does not reflect real-world scenarios where understanding the document’s structure and semantics is crucial (Munikoti et al., 2023; Lai et al., 2024).
-3 The KG-QAGen Benchmark
-KG-QAGen systematically extracts structured question-answer pairs across varying complexity levels for long-context evaluation.
-Using credit agreements (Section 3.1) annotated with a structured label schema (Section 3.2, Appendix A), we construct a knowledge graph (Section 3.3) to enable systematic query construction.
-Questions are designed along three dimensions – multi-hop retrieval, set operations, and answer plurality (Section 3.4) – ensuring scalable, diverse, and automated question extraction (Section 3.5).
-The dataset balances complexity and document coverage, making it a robust benchmark for assessing long-context retrieval.
+RULER (Hsieh et al., 2024) needles partially SEC 10-K & 10-Q 30,000
+BABILong (Kuratov et al., 2024) partially partially Books + Misc 60,000
+MuSiQue (Trivedi et al., 2022) Wikipedia 24,814
+KG-MuLQA-D (ours) SEC Credit Agreements 20,139
+The increasing context length of recent Large Language Models (LLMs) has led to growing interest in evaluating their capabilities.
+Researchers have studied challenges like the Lost in the Middle problem (Liu et al., 2024b), examined the limiting factors of scaling models to long context such as the constraints imposed by RoPE’s base value (Men et al., 2024), and developed techniques to extend context length efficiently (Hooper et al., 2024; Ding et al., 2024).
+Others have focused on improving inference efficiency (Jiang et al., 2024b; Tang et al., 2024) or enhancing long-context utilization and mitigating information loss (Zhang et al., 2024d; Lin et al., 2024; Zhang et al., 2024c).
+Table 1 highlights numerous benchmarks developed to evaluate long-context LLMs on different tasks.
+We introduce the Knowledge-Graph-based Multi-Level Question-Answer Extraction (KG-MuLQA) framework, outlined in Figure 1.
+Our approach uses knowledge-graph document representation to programmatically extract question-answer pairs of varying complexity from gold annotations.
+We define question templates and determine their complexity through three dimensions: the number of hops, the use of set operations, and the plurality of the answer.
+This structured approach ensures scalability, flexibility, and controlled dataset construction, enabling a more precise evaluation of long-context LLMs.
+We apply KG-MuLQA to 170 credit agreements, resulting in KG-MuLQA-D, a dataset of 20,139 QA pairs categorized by five complexity levels.
+We use it to evaluate 16 long-context LLMs and find that even on simple factual queries, some models fail to extract basic information; performance further drops with multi-hop paths and questions involving set operations.
+Our analysis shows that set-based comparisons are especially challenging, with high rates of “Not Found” responses and semantic misinterpretation.
+Our contributions are as follows.
+(1) We present a framework based on knowledge graphs, enabling systematic extraction of QA pairs at varying complexity levels.
+(2) We release a portion of the constructed dataset [^1] https://huggingface.co/datasets/gtfintechlab/KG-MuLQA-D to encourage future work (see Appendix G for details).
+(3) We release our inference pipeline and benchmarking codes [^2] https://github.com/gtfintechlab/KG-MuLQA to facilitate reproducible evaluation of long-context models.
+2 The KG-MuLQA Benchmark
+KG-MuLQA systematically extracts structured QA pairs across varying complexity levels for long-context evaluation.
+Using credit agreements (Section 2.1) annotated with a structured label schema (Section 2.2, Appendix A), we construct document-level knowledge graphs (Section 2.3) to enable systematic query construction.
+Questions are designed along three dimensions – multi-hop reasoning, set operations, and answer plurality (Section 2.4) – ensuring scalable, diverse, and automated question extraction (Section 2.5).
+The dataset balances complexity and document coverage (Appendix A.6), making it a robust long-context LLM benchmark.
 2.1 Credit Agreements
 We sampled 17 credit agreements per year from 2013 to 2022 from SEC EDGAR, totaling 170 documents.
-Table 2 summarizes their length by token and sentence count.
-Their suitability for long-context LLM benchmarking is discussed below, with supporting examples in Appendix A.
-Rich in Entities and Relations
-Credit agreements contain multiple organizations, individuals, roles, and locations, with entities appearing in varied forms throughout the document.
-Some roles have sub-roles (e.g., a lender can be a Swing-Line Lender or Lead Lender).
-Extracting information requires cross-referencing sections, making retrieval a challenging task.
-Ambiguous Co-reference
-Entities are often referenced in multiple ways; for example, “Bank of America, N.A.” may also appear as “Bank of America”.
-Extracting the exact full name can be challenging, as abbreviations or partial mentions may appear throughout the document.
-Structured but not Standardized
-While they follow a general structure (e.g., loan terms, covenants, defaults), variations in phrasing and formatting mean that the same type of information can appear in different sections or be expressed in diverse ways, preventing models from relying on fixed templates and requiring deeper understanding of the document’s content.
-Real-World Significance
-Credit agreements outline multi-million-dollar transactions and define the rights and responsibilities of involved parties.
-Accurate analysis of these documents is essential for financial institutions, legal teams, and regulators to assess risk, ensure compliance, and avoid costly misunderstandings.
-This makes them a highly relevant and impactful target for developing and evaluating QA systems.
+Table 2 includes their length by token and sentence count.
+Their suitability for long-context LLM benchmarking is detailed in Appendix A.1.
 2.2 Annotations
-Table 2: Statistics of the source documents and their structured annotations used to construct KG-QAGen-D.
+Our annotation schema captures key entities and relationships in credit agreements, defining seven label types covering persons, organizations, roles, and locations to represent their structure.
+We outline label definitions in Appendix A.2, placement rules and relationship guidelines in Appendix A.3, while Table 2 includes the frequency statistics for each entity and relation type across documents.
+Each document was independently annotated by 3 annotators.
+Between the second and third (final) rounds of annotating, the inter-annotator agreement reached 80.76%, indicating substantial consistency (see Appendix A.5 for annotation process details).
+2.3 Knowledge Graph Representation
+To enable systematic and scalable QA extraction, we transform annotated credit agreements into knowledge graph representations.
+These graph encode entities of people, organizations, roles, and locations, along with their relationships, in a structured and queryable form.
+Figure 2 illustrates a graph which captures that Jane Smith holds the position of Vice President at ABCDE Inc.
+– an organization that serves as a Lender as well as an Administrative and Documentation Agent, and has its Headquarters at 10 Peachtree....
+By abstracting away document-specific phrasing while preserving these logical links, the graphs support consistent interpretation across agreements and form the basis for creating questions of varying complexity through structured queries.
+Table 2: Statistics of the source documents and their structured annotations used to construct KG-MuLQA-D.
 The table summarizes token and sentence counts, as well as the distribution of annotated entities and relations per document.
@@
 Entities
-Org Name 2 27.71 390
-Org Role 1 13.88 52
+Org Name 2 27.52 390
+Org Role 1 13.82 52
 Org Sub-Role 0 4.26 35
-Person Name 2 12.55 88
-Person Position 2 13.20 88
-Location 1 5.62 168
-Location Type 1 0.43 7
+Person Name 0 12.55 88
+Person Position 0 13.17 88
+Location 0 5.61 168
+Location Type 0 0.45 7
 Relations
@@
 Org Role $\rightarrow$ Org Sub-Role 0 3.74 32
-Org Name $\rightarrow$ Person Name 2 0.97 14
-Person Name $\rightarrow$ Person Position 2 3.04 54
-Org Name $\rightarrow$ Location 1 5.64 178
-Location $\rightarrow$ Location Type 1 0.45 18
-Our annotation schema captures key entities and relationships in credit agreements, defining seven label types covering persons, organizations, roles, and locations to represent their structure.
-Appendix A outlines label definitions, placement rules, and relationship guidelines, while Table 2 illustrates entity relationships and their frequency across documents.
-We annotated the extracted credit agreements using Label-Studio [^1] https://github.com/HumanSignal/label-studio, supporting entity linking and relation extraction for .html files.
-The annotation was carried out by annotators with backgrounds in Computer Science or Mathematics, under the direct supervision of a Master of Finance and a PhD student advised by a Chair Professor of Finance, also a co-author.
-The annotation followed a structured, iterative approach to ensure consistency and resolve edge cases.
-The team first established label definitions and completed an initial annotation round.
-The results were reviewed to identify and discuss edge cases, which led to refinements in the annotation guidelines.
-To enhance reliability, two additional rounds of annotation cleaning were conducted, with annotators reviewing documents they had not previously worked on.
-3.3 Knowledge Graph Representation
+Org Name $\rightarrow$ Person Name 2 12.55 88
+Org Name $\rightarrow$ Person Position 0 13.17 88
+Person Name $\rightarrow$ Person Position 0 12.93 86
+Org Name $\rightarrow$ Location 0 6.32 178
+Location $\rightarrow$ Location Type 0 0.45 18
 Figure 2: Knowledge graph representation of annotated credit agreements.
@@
 All entity names in this figure, including individuals, organizations, and addresses, are fictional and used for illustrative purposes only.
-To enable systematic and scalable QA extraction, we transform annotated credit agreements into a knowledge graph representation.
-This graph encodes entities of people, organizations, roles, and locations, along with their relationships, in a structured and queryable form.
-Figure 2 illustrates a graph which captures that Jane Smith holds the position of Vice President at ABCDE Inc.
-– an organization that serves as a Lender as well as an Administrative and Documentation Agent, and has its Headquarters at 10 Peachtree....
-By abstracting away document-specific phrasing while preserving these logical links, the graph supports consistent interpretation across agreements and forms the basis for creating questions of varying complexity through structured queries.
-We construct the knowledge graph using the Resource Description Framework [^2] https://www.w3.org/TR/rdf11-concepts/ (RDF), which represents data as subject-predicate-object triples, forming a directed labeled graph for knowledge representation.
-Stored in Turtle (*.ttl) format for readability, this structure ensures scalability, and efficient querying across documents using SPARQL [^3] https://www.w3.org/TR/sparql11-query/.
-See Appendix B for knowledge graph construction details.
+We construct knowledge graphs using the Resource Description Framework [^3] https://www.w3.org/TR/rdf11-concepts/ (RDF), which represents data as subject-predicate-object triples, forming a directed labeled graph for knowledge representation.
+Stored in Turtle (*.ttl) format for readability, this structure ensures scalability and efficient querying across documents using SPARQL [^4] https://www.w3.org/TR/sparql11-query/ (see Appendix B for KG construction details).
 2.4 Question Templates and Complexity Dimensions
-We begin by identifying all possible 1-hop QA templates from the knowledge graph, illustrated in Figure 2: questions formed from direct relationships between two entities, such as “What is the position of [Person Name]?" or “Who is the representative of [Org Name]?".
-These base templates are constrained to yield singular answers, forming the simplest form of retrieval.
-It should be noted that not all templates are universally applicable – for instance, the latter question is only valid when the document specifies exactly one representative for the given organization.
-We then systematically expand these templates along three dimensions.
-Number of Hops
-Increasing retrieval depth via intermediate nodes, e.g., “Who is the [Person Position] of the organization located at [Location]?".
-Set Operations
-Introducing comparisons across entities using set logic, e.g., “What are the positions held by [Person Name A] but not [Person Name B] or [Person Name C]?".
-Plurality
-Moving from expecting a singular answer to allowing multiple valid answers, e.g., “What companies are the [(Org Sub-Role +) Org Role] in the agreement?" if multiple roles are held.
-Table 3: This table defines the question templates used to construct KG-QAGen-D, structured along three dimensions: plurality (P), number of hops (H), and set operations (#SO).
+Table 3: This table illustrates the question templates used to construct KG-MuLQA-D, structured along three dimensions: plurality (P), number of hops (H), and set operations (#SO).
 It includes example templates, corresponding knowledge graph query paths, and logical operations involved.
 These dimensions are used to compute the overall complexity level for each QA pair.
-P H #SO Example Template Hop Path Set Operation
+The full list of templates can be found in Appendix C.
+[…]
 0 1 0 What is the position of [Person Name]? Person→Position None
@@
 1 3 2 Who are [Position A]s but not [Position B]s of the company associated with [Location]? Loc→Or→Ps→Pe $A\cap\neg B$
+We begin by identifying all possible 1-hop QA templates from the knowledge graph, illustrated in Figure 2: questions formed from direct relationships between two entities, such as “What is the position of [Person Name]?" or “Who is the representative of [Org Name]?".
+These base templates are constrained to yield singular answers, forming the simplest form of retrieval.
+It should be noted that not all templates are universally applicable – for instance, the latter question is only valid when the document specifies exactly one representative for the given organization.
+We then systematically expand these templates along three dimensions.
+Number of Hops
+Increasing retrieval depth via intermediate nodes, e.g., “Who is the [Person Position] of the organization located at [Location]?".
+Set Operations
+Introducing comparisons across entities using set logic, e.g., “What are the positions held by [Person Name A] but not [Person Name B] or [Person Name C]?".
+Plurality
+Moving from expecting a singular answer to allowing multiple valid answers, e.g., “What companies are the [(Org Sub-Role +) Org Role] in the agreement?".
 2.5 Extraction of QA Pairs from Templates
 Table 3 illustrates how questions are systematically constructed by incrementally increasing complexity across three dimensions introduced in Section 2.4: number of hops (H), plurality (P), and number of set operations (#SO).
-Each row in the table corresponds to a specific question template, with associated values for P, H, and #SO, the query path in the knowledge graph, and the set operation logic used.
+Each row in the table corresponds to a specific question template, with associated values for P, H, and #SO, the query path in a KG, and the set operation logic used.
 Question instances are constructed by dynamically substituting entities into these templates using SPARQL queries over the knowledge graph described in Section 2.3.
@@
 For instance, there are no individuals in an agreement who hold three distinct positions or no organization with the combination of roles required to fulfill a complex query.
-Thus, we limit template depth since such complex conditions are typically not present in the source documents.
-All the created templates can be found in Appendix C.
-To support evaluation, we define a composite question complexity level $L\!=\!H\!+\!P\!+\!\#SO$, and group QA pairs into three categories.
-Easy ($L\!=\!1$): simple, single-hop, singular-answer queries.
-Medium ($42\!\leqslant\!L\!\leqslant\!$ 4): questions with moderate complexity via hops, set operations, or plurality.
-Hard ($L\!=\!5$): questions with the most complex combinations of all three dimensions.
-The resulting dataset, KG-QAGen-D, comprises 20,139 QA pairs from 170 credit agreements, surpassing existing long-context QA benchmarks in size (Table 1).
+Thus, we have to limit template depth since such complex conditions are typically not present in the source documents.
+The list of all the created templates can be found in Appendix C.
+To support evaluation, we define a composite question complexity level L=H+P+#SOL\!=\!H\ \!+\!P\!+\!\#SO, and group QA pairs into three categories.
+Easy (L=1L\!=\!1): simple, single-hop, singular-answer queries.
+Medium ($2\!\leqslant\!L\!\leqslant\!4$): questions with moderate complexity via hops, set operations, or plurality.
+Hard (L=5L\!=\!5): questions with the most complex combinations of all three dimensions.
+The resulting dataset, KG-MuLQA-D, comprises 20,139 QA pairs from 170 credit agreements, surpassing most long-context QA benchmarks in size (Table 1).
 This scale is achieved through efficient question extraction from structured annotations using reusable templates.
@@
 In this section, we describe the setup for our experiments.
-We evaluate multiple long-context LLMs (listed in Section 4.1) across different question complexity levels, as defined in Section 3.4.
+We evaluate multiple long-context LLMs listed in Section 3.1 across different question complexity levels, as defined in Section 2.4.
 Performance is measured using four metrics detailed in Section D.2.
-In addition, we conduct human evaluation of LLM responses on a small sample of QA pairs (see Section D.4).
+In addition, we conduct human evaluation of LLM responses on a small sample of QA pairs (Section D.3) to discover that F1-score and the LLM-as-a-Judge metrics correlate with human’s expectations most.
 Our prompting strategy is outlined in Section D.1.
@@
 3.1 Baselines
-We evaluate 13 proprietary and open-source LLMs on KG-QAGen-D benchmark.
+We evaluate 16 proprietary and open-weight LLMs on KG-MuLQA-D benchmark.
 Proprietary LLMs:
-GPT-4o & GPT-4o-Mini (Achiam et al., 2023; Hurst et al., 2024), Gemini-1.5-Pro & Gemini-1.5-Flash (Team et al., 2023, 2024a), and Gemini-2.0-Flash & Gemini-2.0-Flash-Lite (Team et al., 2024a).
-Open-Source LLMs:
-DeepSeek-V3 (Liu et al., 2024a), Llama-4-Maverick-17B-128E-Instruct-FP8 (AI, 2024a), Llama-4-Scout-17B-16E-Instruct (AI, 2024b), Llama-3.1-405B-Instruct-Turbo & Llama-3.1-70B-Instruct & Llama-3.1-8B-Instruct (Dubey et al., 2024), and Qwen2-72B-Instruct (Yang et al., 2024).
+Gemini-2.0-Flash & Gemini-2.0-Flash-Lite (Team et al., 2024a), Gemini-1.5-Pro & Gemini-1.5-Flash (Team et al., 2023, 2024a), and GPT-4o & GPT-4o-Mini (Achiam et al., 2023; Hurst et al., 2024).
+Open-Weight LLMs:
+DeepSeek-R1 (DeepSeek-AI et al., 2025), DeepSeek-V3 (Liu et al., 2024a), GPT OSS 120B (OpenAI et al., 2025), Llama-4-Maverick-17B-128E-Instruct-FP8 (AI, 2024a), Llama-4-Scout-17B-16E-Instruct (AI, 2024b), Llama-3.1-405B-Instruct-Turbo & Llama-3.1-70B-Instruct & Llama-3.1-8B-Instruct (Dubey et al., 2024), Qwen3-235B-Instruct (Yang et al., 2025), Qwen2-72B-Instruct (Yang et al., 2024).
 Given that our documents average 80k tokens, we restrict our evaluation to models supporting a context length of at least 128k tokens.
 Nonetheless, even long-context models require a carefully designed prompting strategy to extract relevant answers effectively (see Section D.1).
-We made several attempts to benchmark the recently introduced Gemini-2.5-Pro and Gemini-2.5-Flash (Google, 2025) (available in preview mode), but they consistently returned empty responses regardless of configuration.
-We also attempted to evaluate Gemma-2-8k (Team et al., 2024b), Mixtral-Small-32k, and Mixtral-8×22-65k (Jiang et al., 2024a), but these models frequently returned “Not found", making systematic evaluation infeasible.
-For proprietary models and those accessed via Together.AI [^4] https://www.together.ai/, we utilized the LangChain [^5] https://www.langchain.com/ framework, while open-source models were employed using vLLM (Kwon et al., 2023) for efficient inference.
-To ensure reproducibility, we set the temperature to 0.0.
+We made several attempts to evaluate Gemma-2-8k (Team et al., 2024b), Mixtral-Small-32k, and Mixtral-8×22-65k (Jiang et al., 2024a), but these models returned “Not found” in the vast majority of the cases, making systematic evaluation infeasible.
+For proprietary models and those accessed via Together.AI [^5] https://www.together.ai/, we utilized the LangChain [^6] https://www.langchain.com/ framework, while open-weight models were employed using vLLM (Kwon et al., 2023) for efficient inference.
+We set the temperature to 0.0 to support reproducibility of our results.
 3.2 Experiment Setup
+Figure 3: This figure illustrates the multi-stage process used to evaluate long-context LLMs: documents are split into chunks (if exceeding 128K tokens), paired with question batches ($\leqslant$50), and sequentially fed to the model.
+Model responses are collected per chunk and per question batch to ensure evaluation scalability across large documents.
 Our evaluation setup is designed to handle long financial documents and large batches of questions in a way that fits the context length constraints of modern language models.
-We evaluate models by prompting them with a document (or a chunk, if needed) and a batch of 50 associated questions, then measuring the quality of their answers.
-As seen in Figure 3, if a document exceeds the 128K-token context limit, we split it into manageable chunks.
-For each chunk or full document (if it fits), we prompt the model with a batch of 50 questions associated with that segment.
-This design ensures consistent and scalable evaluation across documents of varying lengths.
-See Appendix D.1 for further details.
+As seen in Figure 3, we split a document into chunks if it exceeds 128K tokens: most of the evaluated models have context length of 128K tokens, so it is impossible to evaluate these models without chunking.
+For the larger models, we have to keep the same chunking strategy to ensure a fair comparison between models (see Appendix D.5 for the effect of chunking on evaluation).
+Then, we prompt the model with the document (or chunk) and a batch of 50 questions associated with it: long-context evaluation is very costly, so batching allows at most 3 inference runs per document (given almost 150 questions for some documents).
+This design ensures consistent and scalable evaluation across documents of varying lengths (see Appendix D.1 for further details).
 We use four evaluation metrics: F1 score, Edit Distance, Cosine Similarity, and LLM-as-a-Judge.
-Among these, F1 and LLM-as-a-Judge are highlighted in Table 4 for their complementary strengths – F1 quantifies exact token-level overlap, while LLM-as-a-Judge captures semantic correctness with human-aligned scoring.
-The details are given in Appendix D.2.
-F1 and LLM-as-a-Judge also show the highest correlation with human evaluation metrics (see Appendix D.4).
-Figure 3: This figure illustrates the multi-stage process used to evaluate long-context LLMs: documents are split into chunks (if exceeding 128K tokens), paired with question batches ($\leqslant$50), and sequentially fed to the model.
-Model responses are collected per chunk and per question batch to ensure evaluation scalability across large, real-world documents.
+Among these, F1 and LLM-as-a-Judge are highlighted in Table 4 for their complementary strengths: F1 quantifies exact token-level overlap, while LLM-J captures semantic correctness with human-aligned scoring (see Appendix D.2 for details).
+Both metrics show the highest correlation with human evaluation metrics (see Appendix D.3).
 3.3 Main Results
-The results in Table 4 highlight a significant decline in models performance across all metrics as question complexity increases.
-Llama-3.1-Instruct-Turbo (405B) stands out in the Hard category, achieving an F1 score of 0.3709, significantly higher than all other models, despite not leading in Easy or Medium tasks, which could indicate stronger generalization to difficult tasks.
-However, it was the only model that failed to follow the requested JSON output format, requiring us to develop a separate parser.
-This suggests that as models become larger and more general, their ability to follow specific instructions diminishes.
-Gemini-2.0-Flash (1000K) and Gemini-1.5-Flash (1000K) dominate in Easy and Medium queries but struggle on Hard, indicating a limitation in handling complex retrieval.
-Among open-source models, DeepSeek-V3 (671B) performs consistently but lacks dominance, while Llama-3.1-Instruct (70B) competes well on Medium, surpassing some larger proprietary models.
-Qwen2-Instruct (72B) underperforms across all categories, suggesting size alone does not determine effectiveness.
-Table 4: This table presents the performance of 13 LLMs, evaluated across Easy, Medium, and Hard question categories.
-Metrics include F1 Score and LLM-as-a-Judge rating, capturing both token-level accuracy and semantic correctness.
+As seen in Table 4, the results for simple questions are straightforward to interpret: more advanced models mostly get better scores.
+For medium questions, DeepSeek-V3 outperforming DeepSeek-R1 and Gemini-1.5-Pro outperforming Gemini-2.0 mainly comes from fewer abstentions and “Not Found” cases, which raises recall and F1 but doesn’t reflect stronger reasoning.
+Llama-3.1 405B behaves similarly: it answers more often, boosting both metrics through coverage rather than logic.
+In contrast, Qwen2’s high judge score but low F1 reflects a different pattern: it tends to produce verbose, loosely correct text that seems semantically reasonable, so the judge rates it highly even when exact matches are wrong or incomplete.
+The GPT OSS 120B, however, gets good performance on medium difficulty from its ability to extract entities while minimizing errors, as it strikes a balance between answering and abstaining.
+For hard questions on proprietary models, Gemini-2.0-Flash shows high judged quality but lower F1, while Flash-Lite reverses this trend; Gemini-1.5-Pro remains balanced.
+This difference comes from how often each model answers: more attempts raise F1 through partial overlaps with the gold, while fewer but cleaner attempts earn better judged quality (given responses are not verbose).
+For open-weight models, Llama-4-Maverick achieves a high LLM-J score with moderate F1 for answering selectively with short, factual fragments that look semantically right but miss full token matches.
+Llama-3.1 405B, by contrast, produces longer and more complete answers that align better with references, yielding higher F1 while maintaining strong judged coherence, hence demonstrating the best performance with hard questions.
+Overall, as task complexity increases, models split between answering broadly and answering cautiously.
+This divergence weakens the correlation between F1 and judge scores, since literal overlap and perceived semantic adequacy start reflecting different behaviors.
+4 Analysis
+Table 4: This table presents the performance of 16 LLMs, evaluated across Easy, Medium, and Hard question categories.
+The metrics include the F1 Score and the LLM-as-a-Judge rating, capturing both token-level accuracy and semantic correctness (see the appendix D.2 for extended evaluation).
 The results reveal a consistent decline in performance as question complexity increases, with notable model-specific strengths and weaknesses.
-denotes the models evaluated on a smaller subset due to cost constraints.
+denotes the models evaluated on a smaller subset due to cost constraints (see Appendix D.4 for details).
 Easy Medium Hard
@@
 GPT-4o-Mini 128K 0.4723 2.9690 0.2732 2.0681 0.0274 1.1370
-Open-source LLMs
+Open-weight LLMs
+DeepSeek-R1 685B 128K 0.6269 3.6965 0.3293 2.6354 0.2609 2.0736
 DeepSeek-V3 671B 128K 0.6218 3.5668 0.4796 3.0210 0.0856 1.4795
+GPT OSS 120B 128K 0.5885 3.5839 0.4684 3.1661 0.1136 1.4888
 Llama-4-Maverick-Instruct 400B 1M 0.5552 3.4482 0.3848 2.6976 0.2031 3.1351
@@
 Llama-3.1-Instruct 8B 128K 0.3891 2.6996 0.2962 2.2332 0.1169 1.4315
+Qwen3-Instruct 235B 262K 0.5243 3.2693 0.2760 2.4460 0.1240 1.9578
 Qwen2-Instruct 72B 128K 0.3886 3.3884 0.3100 2.9078 0.0762 1.3288
-5 Analysis
 4.1 Overall Error Trends
-Figure 4: Trends in error types across complexity levels. The template level is given by $P+H+\#SO$.
 As question complexity increases, the LLM’s ability to retrieve and generate correct responses degrades markedly.
-With increase in number of hops or set operations, the percentage distribution of each error type increases (see Figure 4 for more details).
+With increase in number of hops or set operations, the percentage distribution of each error type increases (see Figure 4 for details).
   • Not Found Responses. Increases from 15.00% at Level 1 to 77.39%, reflecting increasing retrieval difficulty.
@@
 For detailed error analysis on examples from each category, see Appendix E, Table 18.
-Misinterpretation of Semantics LLMs frequently misinterpret task-specific terms such as “location type" or “position".
-Rather than grounding responses in document context, models often default to general world knowledge (e.g., answering “city" for a location type) or paraphrase the question, indicating weak comprehension of domain-specific semantics.
+Misinterpretation of Semantics LLMs frequently misinterpret task-specific terms such as “location type” or “position”.
+Rather than grounding responses in document context, models often default to general world knowledge (e.g., answering “city” for a location type), indicating weak comprehension of domain-specific semantics.
 Implicit Information Gaps The model struggles to extract answers that are not explicitly stated but implied through document structure, such as signature blocks or formatting cues.
 For instance, it fails to associate a person with an organization unless the relationship is stated verbatim, ignoring visual or positional indicators that a human reader would readily infer.
-Set Operation Failures To examine the impact of the number of set operations on model performance, we conducted an ablation experiment by varying one dimension within our three-dimensional parameter space while holding the others constant.
+We consider it a potential reason for why models perform slightly better on full documents than in the oracle setting for hard questions (see Appendix D.8 for details).
+Set Operation Failures To examine the impact of the number of set operations on model performance, we conducted an ablation experiment (Figure 11(a)) by varying one dimension within our three-dimensional parameter space while holding the others constant.
 Specifically, we increased the number of set operations while keeping the plurality and hop dimensions fixed.
@@
 The model fails to isolate shared or distinct roles when multiple organizations or individuals are involved, indicating limited capacity for multi-hop or set-based reasoning.
-Long-Context Retrieval Errors Despite the presence of answers in the document, the model often returns “Not found", especially when relevant information appears in the middle of the context.
-This aligns with the “lost-in-the-middle" phenomenon (Liu et al., 2024b), where models disproportionately focus on the beginning or end of long texts, neglecting intermediate content.
+Long-Context Retrieval Errors Despite the presence of answers in the document, models often return “Not Found”, especially when relevant information appears in the middle of the context.
+This aligns with the “lost-in-the-middle” phenomenon (Liu et al., 2024b), where models disproportionately focus on the beginning or end of long texts.
 Our error analysis reveals that while LLMs perform adequately on simple, single-step queries, they falter on multi-hop, set-operation, and implicit-knowledge tasks – particularly those requiring retrieval of details buried in the “middle" of the context.
 Addressing these shortcomings will likely involve a combination of targeted fine-tuning, improved context chunking or retrieval modules, and stronger in-context reasoning prompts.
+5 Related Work
+[…]
+[…]
+Most existing benchmarks do not define explicit question complexity levels, treating all QA instances as uniformly difficult.
+Several recent datasets (Gupta et al., 2024; Zhao et al., 2024c; Zhu et al., 2024; Hsieh et al., 2024; Kuratov et al., 2024) vary task difficulty by changing task type or length, distributing supporting facts in long contexts, or altering context length and needle placement.
+We introduce a granular notion of question complexity, defined along multiple dimensions (multi-hop reasoning, set operations, and answer plurality), allowing systematic control over difficulty and enabling fine-grained error analysis (Section 4.2) that reveals distinct model failure modes.
+While MuSiQue also organizes questions by reasoning depth (2–4 hops), it lacks the multi-dimensional decomposition we propose.
+Our formulation extends beyond hop count to capture qualitatively different reasoning types, which provides a richer understanding of model strengths and weaknesses.
+Figure 4: Trends in error types across complexity levels. The template level is given by P+H+#SOP+H+\#SO.
 6 Discussion & Conclusion
-Our framework integrates structured annotations, knowledge graph construction, and templated QA extraction (Sections 3.2-3.5) to produce a dataset with finely controlled question complexity over real-world long document.
-This allows systematic evaluation across various dimensions with identifying nuanced model failures.
-Despite high scores on simple factual questions, even the strongest LLMs struggle with compositional logical inference, especially set-based comparisons and multi-hop paths, frequently failing to interpret implicit relationships and falling back on world knowledge instead of grounding answers in context (Section 5).
+Our work introduces KG-MuLQA, a framework that programmatically constructs long-context QA benchmarks using knowledge-graph representations of real-world documents.
+We begin by annotating credit agreements with a structured schema to capture entities and relations, convert them into RDF-based knowledge graphs, and generate multi-level QA pairs via deterministic SPARQL templates that vary in hop count, set operations, and answer plurality.
+We then evaluate 16 long-context LLMs across these controlled complexity levels using 4 complementary metrics and perform human evaluation to validate metric alignment.
+Our quantitative and qualitative analyses reveal consistent weaknesses in multi-hop retrieval, set-based reasoning, and implicit relation grounding, establishing KG-MuLQA as a scalable and interpretable framework for assessing long-context comprehension.
+Human Performance Evaluation
+Appendix D.7 shows human performance remains stable across difficulty.
+This baseline both grounds our metric choices and indicates model errors stem from reasoning/retrieval limits rather than dataset artifacts.
 Applicability and Generalizability
-As discussed in Sections 1 and 3.1, the ability to process long documents is essential for real-world QA tasks.
+As discussed in Appendix A.1, the ability to process long documents is essential for real-world QA tasks.
 Our benchmark illustrates this challenge using credit agreements, where information is densely structured, scattered, and cross-referenced.
-However, legal contracts, policy documents, patent filings, and other non-financial documents all exhibit similarly complex entity structures.
-Even in non-legal domains, such as medical records or scientific articles, multi-hop relationships and compositional multi-hop inference over entities are common, making our methodology broadly applicable.
-Opportunities and Extensions
-While our benchmark focuses on pre-trained LLMs, the structured design of KG-QAGen enables future adaptation, such as domain-specific fine-tuning.
-Expanding the dataset beyond U.S.-based financial documents could support geographically diverse evaluation.
+Legal contracts, policy documents, patent filings, and other non-legal such as medical records or scientific articles all exhibit similarly complex entity structures, making our methodology broadly applicable.
+To illustrate the applicability and generalizability of our framework, we performed a full-scale ablation study (see Appendix F) on medical documents, including defining the annotation schema, constructing question templates, etc.
+Limitations
+The created QA pairs are based on financial credit agreements only.
+As with most existing QA benchmarks, the benchmark favors deductive logic because the dataset includes graph-representable questions.
+Ethical Considerations
+We generally avoid mentioning specific agreements and their details.
+Instead, we use placeholders such as ABCDE Inc.
+for organization names, Richard Roe for person names, and 10 Peachtree...
+for locations to illustrate the structure of agreement entities and relations.
+However, in our annotation guide (Appendix A) and error analysis (Appendix E), we include screenshots of real document annotations and mention specific examples to highlight the complexity of these agreements.
+Moreover, we store the raw HTML files of the agreements in our GitHub repository, enabling other researchers to reproduce our benchmark or conduct further experiments.
+As these documents are publicly available through the SEC’s EDGAR system, their inclusion for both illustrative and experimental purposes raises no ethical concerns.
 Appendix A Annotation Guide
-A.1 Labels
+A.1 Suitability of Credit Agreements
+Credit agreements are particularly well-suited for long-context evaluation because they are rich in entities and relations, involving numerous organizations, individuals, roles, and locations that appear in varied forms throughout the document, making both retrieval and reasoning challenging.
+They also exhibit ambiguous co-reference, as entities may be referred to inconsistently (e.g., “Bank of America, N.A.” vs.
+“Bank of America”), demanding precise resolution of abbreviations and partial mentions.
+While structured but not standardized, these documents follow broad templates (e.g., loan terms, covenants, defaults) but vary greatly in phrasing and layout, preventing reliance on fixed patterns and requiring genuine comprehension.
+Finally, their real-world significance as legally binding records of multi-million-dollar transactions makes accurate understanding essential for assessing risk, ensuring compliance, and supporting financial and regulatory decisions.
+Supporting examples illustrating these aspects are provided in the latter subsections of this Appendix.
+A.2 Labels
 Our annotation framework captures key entities and relationships within credit agreements, focusing on seven labels that define organizational structure, roles, and locations.
 Organization Name
-The name of a company or financial institution mentioned in the agreement. Examples: "Goldman Sachs", "JP Morgan Chase".
+The name of a company or financial institution mentioned in the agreement. Examples: “Goldman Sachs”, “JP Morgan Chase”.
 Organization Role
-The role that an organization holds in the agreement. Examples: "Lender", "Borrower", "Guarantor".
+The role that an organization holds in the agreement. Examples: “Lender”, “Borrower”, “Guarantor”.
 Organization Sub-Role
-A more specific classification within an Organization Role that provides additional detail. Examples: "Lead Lender", "Administrative Agent".
+A more specific classification within an Organization Role that provides additional detail. Examples: “Lead Lender”, “Administrative Agent”.
 Person Name
-The name of an individual explicitly mentioned in the agreement. Examples: "John Doe", "Jane Smith".
+The name of an individual explicitly mentioned in the agreement. Examples: “John Doe”, “Jane Smith”.
 Person Position
-The title or role held by a person within an organization. Examples: "Vice President", "Loan Officer".
+The title or role held by a person within an organization. Examples: “Vice President”, “Loan Officer”.
 Location
-A physical address associated with an organization. Examples: "35 Peachtree St, Atlanta, GA".
+A physical address associated with an organization. Examples: “35 Peachtree St, Atlanta, GA”.
 Location Type
-A classification of a Location, providing additional information on its function. Example: "Headquarters", "Branch Office".
-A.2 General rules
+A classification of a Location, providing additional information on its function. Example: “Headquarters”, “Branch Office”.
+A.3 General rules
 Unique Entity Annotation
@@
 Organization Roles, Organization Sub-Roles, Person Positions, and Location Types are annotated once for each unique entity they describe.
-Example: In Figure 5(a), "BANK OF AMERICA, N.A." appears multiple times but is annotated only once.
-However, its roles, "Administrative Agent" and "Lender", are each annotated where they are explicitly stated.
+Example: In Figure 5(a), “BANK OF AMERICA, N.A.” appears multiple times but is annotated only once.
+However, its roles, “Administrative Agent” and “Lender”, are each annotated where they are explicitly stated.
 If another organization is a lender and/or an administrative agent, those role labels will be annotated as close as possible to the name label.
-Similarly for "Alexandra M.
-Knights" and her position "Authorized Signatory".
+Similarly for “Alexandra M.
+Knights” and her position “Authorized Signatory”.
 Ordered Search for Annotations
@@
 Relation Annotations
-  • Org Name $\rightarrow$ Org Role: Each organization name is annotated once per document, while its role is annotated as close to the name as possible.
-If multiple organizations share the same role (e.g., Borrower) and it is mentioned collectively in the signature box, the role is annotated once and linked to all relevant organizations.
-Otherwise, each occurrence of a role is annotated separately and linked to the corresponding organization.
-  • Org Role $\rightarrow$ Org Sub-Role: If an organization has a specific sub-role (e.g., Parent Borrower), its role is annotated separately from others, and the sub-role is linked to it.
-The connection follows a hierarchical structure, linking the organization to its role, then the role to its sub-role.
-  • Org Name $\rightarrow$ Person Name: Each person’s name is annotated once per document and linked to a single organization (see A.3 for special scenarios).
-  • Person Name $\rightarrow$ Person Position: A person’s position is annotated as close to their name as possible.
-Individuals may hold multiple positions, which are all linked to the same person annotation.
-In some cases, position mentions may be missing (see Figure 5(d) and A.3).
-  • Org Name $\rightarrow$ Location: Locations are annotated once per document and linked to the relevant organization.
-  • Location $\rightarrow$ Location Type: Each location has exactly one type, which is annotated as close to the location name as possible.
+[…]
+[…]
+[…]
+[…]
+[…]
+  • $\rightarrow Location$ Type: Each location has exactly one type, which is annotated as close to the location name as possible.
 Location types are often missing, and in such cases, no annotation is applied (see Fig 5(d)).
@@
 Since our annotation structure does not include relations between labels of the same type, these links are easily identifiable.
-During post-processing, entities connected by continuation links are automatically merged with spaces, ensuring a single coherent label.
-A.3 Main Edge Cases
+During post-processing, entities connected by continuation links are automatically merged with spaces into a single coherent label.
+A.4 Main Edge Cases
 Parent Borrower and Related Entities
@@
 Smallest Document
-One of the extracted credit agreements has only 2 organizations involved, and their roles are mentioned neither in the signature box nor in the header (Figure 5(f)).
-Moreover, the role of the lender is not mentioned anywhere in the document, while the role of the borrower is mentioned once (Figure 5(g)).
+One of the extracted credit agreements involves only 2 organizations, and their roles are mentioned neither in the signature box nor in the header (Figure 5(f)).
+Moreover, the role of the lender is not mentioned in the document, while the role of the borrower is mentioned once (Figure 5(g)).
+A.5 Annotation Process Details
+We annotated the extracted credit agreements using Label-Studio [^7] https://github.com/HumanSignal/label-studio, supporting entity linking and relation extraction for .html files.
+The annotation was carried out by annotators with backgrounds in Computer Science or Mathematics, under the direct supervision of a Master of Finance and a PhD student advised by a Chair Professor of Finance, also a co-author.
+The annotation followed a structured, iterative approach to ensure consistency and resolve edge cases.
+The team first established label definitions and completed an initial annotation round.
+The results were reviewed to identify and discuss edge cases (Appendix A.4), which led to refinements in the annotation guidelines and the second round of annotations (inter-annotator agreement was not calculated at this point, as the guidelines were evolving).
+Following the second annotation round, we constructed knowledge-graph (KG) document representations, sampled graphs, and compared them against the ground-truth annotations.
+Insights from this qualitative assessment (Appendix B.1) informed a third annotation round, after which we applied an additional cleaning script to the KG representations.
+A.6 Complexity vs coverage
+KG-MuLQA defines seven entity and six relation types, and this choice reflects a deliberate trade-off between semantic coverage and compositional complexity.
+Expanding the ontology would increase topical breadth but would not necessarily yield deeper reasoning requirements.
+Our goal is to evaluate long-context comprehension under settings where information is distributed across distant parts of a document and must be retrieved, composed, and reasoned over.
+Thus, the benchmark emphasizes complex inference chains (multi-hop reasoning, set operations, and answer plurality) over the schema coverage.
+Answering questions from our dataset requires linking dispersed mentions, disambiguating roles, and integrating partial evidence: operations that jointly test both retrieval and reasoning.
+This structure allows us to probe failure modes of long-context models in a controlled environment, balancing interpretability, scalability, and difficulty without relying on full document coverage.
 (a) General and most common annotation in the signature box.
@@
 Figure 6: SPARQL Query Example for Structured QA Extraction.
+B.1 Qualitative Analysis of the Constructed KGs
+After generating question-answer pairs from the extracted graphs, we observed two common issues affecting the quality and consistency of the QAs.
+  • Duplicate entries: Multiple questions or answers representing the same entity or relation, differing only by capitalization, whitespace, punctuation, or pluralization.
+  • Noisy text: URL-encoded characters (e.g., %26, %20) and miscellaneous special characters were present in questions and answers.
+To address these issues, we applied a structured cleaning procedure after QA generation, ensuring the graphs themselves were not altered but the extracted QAs were reliable:
+  • Text normalization:
+    • Convert all text to lowercase.
+    • Remove URL-encoded sequences and special characters, preserving commas for multiple answers.
+    • Collapse multiple whitespace characters.
+  • Duplicate answer removal:
+    • Split multi-answer entries by commas and clean each entry.
+    • Normalize text (lowercase, remove spaces) and reduce plural forms to singular where appropriate.
+    • Remove duplicates while preserving the first occurrence.
+    • Log all removed duplicates, including document identifiers, for traceability.
 Appendix C All templates
-Table 5: Level 1 Question Templates.
-Template Based On P H #SO
-What is the position of [Person Name]? [if one] 0 1 0
-In what organization does [Person Name] work? 0 1 0
-Who is the representative of [Org Name]? [if one] 0 1 0
-What is the role of [Org Name] in the agreement? [if one] 0 1 0
-What company is the [Org Role (+ Sub-Role)] in the agreement? [if one] 0 1 0
-What is the location of [Org Name]? [if one] 0 1 0
-Which company is associated with [Location]? 0 1 0
-What type of location is [Location] (e.g., Headquarters, Trade Operations, etc.)? [if one] 0 1 0
-Table 6: Level 2 Question Templates.
-Template Based On P H #SO
-What are the positions of [Person Name]? What is the position of [Person Name]? [if one] 1 1 0
-Who are the representatives of [Org Name]? Who is the representative of [Org Name]? [if one] 1 1 0
-What are the roles of [Org Name] in the agreement? What is the role of [Org Name] in the agreement? [if one] 1 1 0
-What companies are the [Org Role (+ Sub-Role)] in the agreement? What company is the [Org Role (+ Sub-Role)] in the agreement? [if one] 1 1 0
-What are the locations of [Org Name]? What is the location of [Org Name]? [if one] 1 1 0
-What types of location is [Location] (e.g., Headquarters, Trade Operations, etc.)? What type of location is [Location] (e.g., Headquarters, Trade Operations, etc.)? [if one] 1 1 0
-What is the position held by both [Person Name 1] and [Person Name 2]? [if one] What is the position of [Person Name]? [if one] 0 1 1
-What role do both [Org Name 1] and [Org Name 2] have in the agreement? [if one] What is the role of [Org Name] in the agreement? [if one] 0 1 1
-What company is both the [Org Role (+ Sub-Role) 1] and [Org Role (+ Sub-Role) 2] in the agreement? [if one] What company is the [Org Role (+ Sub-Role)] in the agreement? [if one] 0 1 1
-Who is the [Person Position] of [Org Name]? [if one] 0 2 0
-What is the role in the agreement of the company where [Person Name] is employed? [if one] What is the role of [Org Name] in the agreement? [if one] 0 2 0
-What is the role in the agreement of the company associated with [Location]? [if one] What is the role of [Org Name] in the agreement? [if one] 0 2 0
-What is the [Location Type] office of [Org Name]? 0 2 0
-Table 7: Level 3 Question Templates.
-Template Based On P H #SO
-What are the positions held by both [Person Name 1] and [Person Name 2]? What are the positions of [Person Name]? What is the position held by both [Person Name 1] and [Person Name 2]? [if one] 1 1 1
-What roles do both [Org Name 1] and [Org Name 2] have in the agreement?
-What are the roles of [Org Name] in the agreement?
-What role do both [Org Name 1] and [Org Name 2] have in the agreement?
-[if one] 1 1 1
-What companies are both the [Org Role (+ Sub-Role) 1] and [Org Role (+ Sub-Role) 2] in the agreement? What companies are the [Org Role (+ Sub-Role)] in the agreement? 1 1 1
-What is the position held by [Person Name 1] but not by [Person Name 2]? [if one] What is the position held by both [Person Name 1] and [Person Name 2]? [if one] 0 1 2
-What role does [Org Name 1] have in the agreement which is not the role of [Org Name 2]? [if one] What role do both [Org Name 1] and [Org Name 2] have in the agreement? [if one] 0 1 2
+This appendix contains the defined question templates.
+Table 5: Level 1 Question Templates with Counts.
+Template Based On P H #SO Count
+What is the position of [Person Name]? [if one] 0 1 0 1252
+In what organization does [Person Name] work? 0 1 0 1288
+Who is the representative of [Org Name]? [if one] 0 1 0 1358
+What is the role of [Org Name] in the agreement? [if one] 0 1 0 845
+What company is the [Org Role (+ Sub-Role)] in the agreement? [if one] 0 1 0 444
+What is the location of [Org Name]? [if one] 0 1 0 790
+Which company is associated with [Location]? 0 1 0 530
+What type of location is [Location] (e.g., Headquarters, Trade Operations, etc.)? [if one] 0 1 0 43
+Table 6: Level 2 Question Templates with Counts.
+Template Based On P H #SO Count
+Who is the [Person Position] of [Org Name]? [if one] 0 2 0 2274
+What is the position held by both [Person Name 1] and [Person Name 2]? [if one] What is the position of [Person Name]? [if one] 0 1 1 1612
+What is the role in the agreement of the company where [Person Name] is employed? [if one] What is the role of [Org Name] in the agreement? [if one] 0 2 0 613
+What are the roles of [Org Name] in the agreement? What is the role of [Org Name] in the agreement? [if one] 1 1 0 409
+What companies are the [Org Role (+ Sub-Role)] in the agreement? What company is the [Org Role (+ Sub-Role)] in the agreement? [if one] 1 1 0 304
+What role do both [Org Name 1] and [Org Name 2] have in the agreement? [if one] What is the role of [Org Name] in the agreement? [if one] 0 1 1 257
+What is the role in the agreement of the company associated with [Location]? [if one] What is the role of [Org Name] in the agreement? [if one] 0 2 0 180
+Who are the representatives of [Org Name]? Who is the representative of [Org Name]? [if one] 1 1 0 160
+What are the positions of [Person Name]? What is the position of [Person Name]? [if one] 1 1 0 136
+What are the locations of [Org Name]? What is the location of [Org Name]? [if one] 1 1 0 91
+In what organizations does [Person Name] work? In what organization does [Person Name] work? 1 1 0 90
+What is the [Location Type] office of [Org Name]? 0 2 0 70
+What types of location is [Location] (e.g., Headquarters, Trade Operations, etc.)? What type of location is [Location] (e.g., Headquarters, Trade Operations, etc.)? [if one] 1 1 0 11
+Table 7: Level 3 Question Templates with Counts.
+Template Based On P H #SO Count
+What are the positions held by both [Person Name 1] and [Person Name 2]? What are the positions of [Person Name]? 1 1 1 3
+What is the position held by [Person Name 1] but not by [Person Name 2]? [if one] What are the positions held by both [Person Name 1] and [Person Name 2]? 0 1 2 8
+What roles do both [Org Name 1] and [Org Name 2] have in the agreement? What are the roles of [Org Name] in the agreement? 1 1 1 92
+What role does [Org Name 1] have in the agreement which is not the role of [Org Name 2]? [if one] What roles do both [Org Name 1] and [Org Name 2] have in the agreement? 0 1 2 278
 What company is the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] in the agreement?
-[if one] What company is both the [Org Role (+ Sub-Role) 1] and [Org Role (+ Sub-Role) 2] in the agreement?
-[if one] 0 1 2
-Who are the [Person Position]s of [Org Name]? Who is the [Person Position] of [Org Name]? [if one] 1 2 0
-What are the roles in the agreement of the company where [Person Name] is employed? What is the role in the agreement of the company where [Person Name] is employed? [if one] 1 2 0
-What are the roles in the agreement of the company associated with [Location]? What is the role in the agreement of the company associated with [Location]? [if one] 1 2 0
-Who is the [Person Position 1] and [Person Position 2] of [Org Name]? [if one] Who is the [Person Position] of [Org Name]? [if one] 0 2 1
+[if one] What companies are both the [Org Role (+ Sub-Role) 1] and [Org Role (+ Sub-Role) 2] in the agreement?
+0 1 2 56
+Who are the [Person Position]s of [Org Name]? Who is the [Person Position] of [Org Name]? [if one] 1 2 0 57
+Who is the [Person Position 1] and [Person Position 2] of [Org Name]? [if one] Who is the [Person Position] of [Org Name]? [if one] 0 2 1 1346
+What are the roles in the agreement of the company where [Person Name] is employed? What is the role in the agreement of the company where [Person Name] is employed? [if one] 1 2 0 318
+What are the roles in the agreement of the company associated with [Location]? What is the role in the agreement of the company associated with [Location]? [if one] 1 2 0 169
 Who is the [Person Position] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement?
 [if one, and the company should be uniquely identifiable] Who is the [Person Position] of [Org Name]?
-[if one] 0 3 0
-Who is the [Person Position] of the company associated with [Location]? [if one] Who is the [Person Position] of [Org Name]? [if one] 0 3 0
-Who is the [Person Position] of the company associated where [Person Name] is employed? [if one] Who is the [Person Position] of [Org Name]? [if one] 0 3 0
+[if one] 0 3 0 262
+Who is the [Person Position] of the company associated with [Location]? [if one] Who is the [Person Position] of [Org Name]? [if one] 0 3 0 616
+Who is the [Person Position] of the company where [Person Name] is employed? [if one] Who is the [Person Position] of [Org Name]? [if one] 0 3 0 289
 What is the address of [Location Type] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement?
 [if one, and the company should be uniquely identifiable] What is the [Location Type] office of [Org Name]?
-0 3 0
-What is the address of [Location Type] of the company associated with [Location (different)]? What is the [Location Type] office of [Org Name]? 0 3 0
-What is the address of [Location Type] of the company where [Person Name] is employed? What is the [Location Type] office of [Org Name]? 0 3 0
+0 3 0 13
+What is the address of [Location Type] of the company where [Person Name] is employed? What is the [Location Type] office of [Org Name]? 0 3 0 75
+Who is the [Person Position] of the company where [Person Name] is employed? [if one] Who is the [Person Position] of [Org Name]? [if one] 0 3 0 289
 Table 8: Level 4 Question Templates, part 1.
-Template Based On P H #SO
+Template Based On P H #SO Count
 What are the positions held by [Person Name 1] but not by [Person Name 2]?
@@
 What is the position held by [Person Name 1] but not by [Person Name 2]?
-[if one] 1 1 2
+[if one] 1 1 2 2
 What roles does [Org Name 1] have in the agreement which are not the roles of [Org Name 2]?
@@
 What role does [Org Name 1] have in the agreement which is not the role of [Org Name 2]?
-[if one] 1 1 2
+[if one] 1 1 2 134
 What companies are the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] in the agreement?
@@
 What company is the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] in the agreement?
-[if one] 1 1 2
-What is the position held by [Person Name 1] but not by [Person Name 2] or [Person Name 3]? [if one] What is the position held by [Person Name 1] but not by [Person Name 2]? [if one] 0 1 3
-What is the position held by [Person Name 1] and [Person Name 2] but not by [Person Name 3]? [if one] What is the position held by [Person Name 1] but not by [Person Name 2]? [if one] 0 1 3
+[if one] 1 1 2 124
+What is the position held by [Person Name 1] but not by [Person Name 2] or [Person Name 3]? [if one] What is the position held by [Person Name 1] but not by [Person Name 2]? [if one] 0 1 3 50
+What is the position held by [Person Name 1] and [Person Name 2] but not by [Person Name 3]? [if one] What is the position held by [Person Name 1] but not by [Person Name 2]? [if one] 0 1 3 4
 What role do [Org Name 1] and [Org Name 2] have in the agreement which is not the role of [Org Name 3]?
 [if one] What role does [Org Name 1] have in the agreement which is not the role of [Org Name 2]?
-[if one] 0 1 3
+[if one] 0 1 3 0
 What role does [Org Name 1] have in the agreement which is not the role of [Org Name 2] or [Org Name 3]?
 [if one] What role does [Org Name 1] have in the agreement which is not the role of [Org Name 2]?
-[if one] 0 1 3
+[if one] 0 1 3 0
 What company is the [Org Role (+ Sub-Role) 1] and [Org Role (+ Sub-Role) 2] but not the [Org Role (+ Sub-Role) 3] in the agreement?
 [if one] What company is the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] in the agreement?
-[if one] 0 1 3
+[if one] 0 1 3 1553
 What company is the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] [Org Role (+ Sub-Role) 3] in the agreement?
 [if one] What company is the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] in the agreement?
-[if one] 0 1 3
+[if one] 0 1 3 816
 Table 9: Level 4 Question Templates, part 2.
-Template Based On P H #SO
+Template Based On P H #SO Count
 Who are the both [Person Position 1]s and [Person Position 2]s of [Org Name]?
@@
 Who is the [Person Position 1] and [Person Position 2] of [Org Name]?
-[if one] 1 2 1
+[if one] 1 2 1 0
 Who are the [Person Position]s of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement?
 [the company should be uniquely identifiable] Who is the [Person Position] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement?
-[if one, and the company should be uniquely identifiable] 1 3 0
-Who are the [Person Position]s of the company associated with [Location]? Who is the [Person Position] of the company associated with [Location]? [if one] 1 3 0
-Who are the [Person Position]s of the company associated where [Person Name] is employed? Who is the [Person Position] of the company associated where [Person Name] is employed? [if one] 1 3 0
+[if one, and the company should be uniquely identifiable] 1 3 0 40
+Who are the [Person Position]s of the company associated with [Location]? Who is the [Person Position] of the company associated with [Location]? [if one] 1 3 0 22
+Who are the [Person Position]s of the company associated where [Person Name] is employed? Who is the [Person Position] of the company associated where [Person Name] is employed? [if one] 1 3 0 5
 Who is the [Person Position] of the company which is both the [Org Role(-s) (+ Sub-Role(-s)) 1] and the [Org Role(-s) (+ Sub-Role(-s)) 2] in the agreement?
 [if one, and the company should be uniquely identifiable] Who is the [Person Position] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement?
-[if one, and the company should be uniquely identifiable] 0 3 1
+[if one, and the company should be uniquely identifiable] 0 3 1 187
 Who is both the [Person Position 1] and [Person Position 2] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement?
 [if one, and the company should be uniquely identifiable] Who is the [Person Position] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement?
-[if one, and the company should be uniquely identifiable] 0 3 1
-Who is both the [Person Position 1] and [Person Position 2] of the company associated with [Location]? [if one] Who is the [Person Position] of the company associated with [Location]? [if one] 0 3 1
+[if one, and the company should be uniquely identifiable] 0 3 1 61
+Who is both the [Person Position 1] and [Person Position 2] of the company associated with [Location]?
+[if one] Who is the [Person Position] of the company associated with [Location]?
+[if one] 0 3 1 59
 Who is both the [Person Position 1] and [Person Position 2] of the company associated where [Person Name] is employed?
 [if one] Who is the [Person Position] of the company associated where [Person Name] is employed?
-[if one] 0 3 1
+[if one] 0 3 1 23
 What is the address of the [Location Type] office of the company which is both the [Org Role(-s) (+ Sub-Role(-s)) 1] and the [Org Role(-s) (+ Sub-Role(-s)) 2] in the agreement?
 [if one, and the company should be uniquely identifiable] What is the [Location Type] office of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement?
-[if one, and the company should be uniquely identifiable] 0 3 1
+[if one, and the company should be uniquely identifiable] 0 3 1 14
 Table 10: Level 5 Question Templates, part 1.
-Template Based On P H #SO
+Template Based On P H #SO Count
 What are the positions held by [Person Name 1] but not by [Person Name 2] or [Person Name 3]?
@@
 What is the position held by [Person Name 1] but not by [Person Name 2] or [Person Name 3]?
-[if one] 1 1 3
+[if one] 1 1 3 8
 What are the positions held by [Person Name 1] and [Person Name 2] but not by [Person Name 3]?
@@
 What is the position held by [Person Name 1] and [Person Name 2] but not by [Person Name 3]?
-[if one] 1 1 3
+[if one] 1 1 3 49
 What roles do [Org Name 1] and [Org Name 2] have in the agreement which are not the roles of [Org Name 3]?
@@
 What role do [Org Name 1] and [Org Name 2] have in the agreement which is not the role of [Org Name 3]?
-[if one] 1 1 3
+[if one] 1 1 3 0
 What roles does [Org Name 1] have in the agreement which are not the roles of [Org Name 2] or [Org Name 3]?
@@
 What role does [Org Name 1] have in the agreement which is not the role of [Org Name 2] or [Org Name 3]?
-[if one] 1 1 3
+[if one] 1 1 3 0
 What companies are the [Org Role (+ Sub-Role) 1] and [Org Role (+ Sub-Role) 2] but not the [Org Role (+ Sub-Role) 3] in the agreement?
@@
 What company is the [Org Role (+ Sub-Role) 1] and [Org Role (+ Sub-Role) 2] but not the [Org Role (+ Sub-Role) 3] in the agreement?
-[if one] 1 1 3
+[if one] 1 1 3 122
 What companies are the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] [Org Role (+ Sub-Role) 3] in the agreement?
@@
 What company is the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] [Org Role (+ Sub-Role) 3] in the agreement?
-[if one] 1 1 3
+[if one] 1 1 3 0
 Table 11: Level 5 Question Templates, part 2.
-Template Based On P H #SO
+Template Based On P H #SO Count
 Who are the [Person Position]s of the company which is both the [Org Role(-s) (+ Sub-Role(-s)) 1] and the [Org Role(-s) (+ Sub-Role(-s)) 2] in the agreement?
@@
 [the company should be uniquely identifiable] Who is the [Person Position] of the company which is both the [Org Role(-s) (+ Sub-Role(-s)) 1] and the [Org Role(-s) (+ Sub-Role(-s)) 2] in the agreement?
-[if one, and the company should be uniquely identifiable] 1 3 1
+[if one, and the company should be uniquely identifiable] 1 3 1 0
 Who are both the [Person Position 1] and [Person Position 2] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement?
@@
 [the company should be uniquely identifiable] Who is both the [Person Position 1] and [Person Position 2] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement?
-[if one, and the company should be uniquely identifiable] 1 3 1
+[if one, and the company should be uniquely identifiable] 1 3 1 0
 Who are both the [Person Position 1] and [Person Position 2] of the company associated with [Location]?
@@
 Who is both the [Person Position 1] and [Person Position 2] of the company associated with [Location]?
-[if one] 1 3 1
+[if one] 1 3 1 0
 Who are both the [Person Position 1] and [Person Position 2] of the company associated where [Person Name] is employed?
@@
 Who is both the [Person Position 1] and [Person Position 2] of the company associated where [Person Name] is employed?
-[if one] 1 3 1
+[if one] 1 3 1 0
 Who is the [Person Position 1] but not [Person Position 2] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement?
 [if one, and the company should be uniquely identifiable] Who is both the [Person Position 1] and [Person Position 2] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement?
-[if one, and the company should be uniquely identifiable] 0 3 2
+[if one, and the company should be uniquely identifiable] 0 3 2 2
 Who is the [Person Position 1] but not [Person Position 2] of the company associated with [Location]?
 [if one] Who is both the [Person Position 1] and [Person Position 2] of the company associated with [Location]?
-[if one] 0 3 2
+[if one] 0 3 2 1
 Who is the [Person Position 1] but not [Person Position 2] of the company associated where [Person Name] is employed?
 [if one] Who is both the [Person Position 1] and [Person Position 2] of the company associated where [Person Name] is employed?
-[if one] 0 3 2
+[if one] 0 3 2 3
+Who is the [Person Position] of the company which is the [Org Role(-s) (+ Sub-Role(-s)) 1] but not the [Org Role(-s) (+ Sub-Role(-s)) 2] in the agreement?
+[if one, and the company should be uniquely identifiable] Who is the [Person Position] of the company which is both the [Org Role(-s) (+ Sub-Role(-s)) 1] and the [Org Role(-s) (+ Sub-Role(-s)) 2] in the agreement?
+[if one, and the company should be uniquely identifiable] 0 3 2 508
 What is the [Location Type] office of the company which is both the [Org Role(-s) (+ Sub-Role(-s)) 1] but not the [Org Role(-s) (+ Sub-Role(-s)) 2] in the agreement?
 [if one, and the company should be uniquely identifiable] What is the [Location Type] office of the company which is both the [Org Role(-s) (+ Sub-Role(-s)) 1] and the [Org Role(-s) (+ Sub-Role(-s)) 2] in the agreement?
-[if one, and the company should be uniquely identifiable] 0 3 2
+[if one, and the company should be uniquely identifiable] 0 3 2 13
 Appendix D Evaluation Details
 D.1 Prompting Strategy
-Batch QA Prompt [SYSTEM INPUT]
-You are a financial expert, and your task is to answer the questions given to you in batches about the provided credit agreement.
+We implement a two-stage prompting pipeline for QA inference over long documents.
+In the first stage, the document is either processed in full or split into chunks (depending on the context limit of the LLM), and each chunk is paired with a batch of up to 50 questions.
+We then apply a one-shot, instruction-based prompt (Figure 7(a)) that instructs the LLM to act as a financial expert and return an array of answers – either grounded in the document or explicitly stating “Not found" – in valid JSON format.
+The effect of chunk-based evaluation is discussed in Appendix D.5.
+The expected format is enforced both through the prompt instructions and through structured parsing logic.
+The parser first attempts to extract a JSON object from the model’s response, supporting both raw JSON and code-fenced output.
+If parsing fails, the system falls back to extracting answers from list-formatted outputs or reattempts the query, up to a maximum number of retries.
+All models are queried with temperature=0.0.
+In the second stage, for questions where answers may be distributed across multiple chunks, a merging prompt is issued (Figure 7(b)).
+This prompt consolidates partial answers from earlier responses and asks the model to output a single final answer per question, again using the same strict JSON format.
+This strategy allows the model to incorporate information across multiple views of the document.
+Regarding resources, as mentioned in Section 3.1, for open-weight models we are using Together.AI for some of the models. For the other open-weight models, we used 1 H200 GPU.
+D.2 Evaluation Metrics
+We employ four complementary metrics to evaluate model responses (Table 12).
+  • Word-Level F1 Score  Token-level precision and recall against the gold answer, ensures strict factual correctness in question answering.
+  • Normalized Levenshtein Distance  Character-level string distance, indicating surface-level divergence (described by Yujian and Bo (2007)).
+  • Cosine Similarity  Evaluates the semantic alignment, with robustness to minor changes.
+  • LLM-as-a-Judge  Gemini-2.0-Pro is prompted to score each prediction from 1 to 5 based on its semantic correctness, given the question and reference answer.
+We exclude ROUGE, BLEU, and similar metrics because they are known to perform poorly on short factual answers, tend to over-penalize surface-level variations and are not well-aligned with human judgment in our setting (Blagec et al., 2022; Chaganty et al., 2018).
+D.3 Human Evaluation Metrics
+We conducted a human evaluation of the LLM-generated responses using three independent evaluators.
+Two of these evaluators were annotators familiar with credit agreements, while the third evaluator had no prior exposure to the documents.
+Each evaluator was requested to rate the LLM responses on a scale from 1 to 5, according to the criteria below.
+A total of 100 questions were sampled from multiple templates to ensure maximal diversity in the human evaluation task.
+  • 1   response is irrelevant or does not match the actual answer;
+  • 2   response is somewhat relevant but does not match the actual answer;
+  • 3   response shows slight similarity but lacks accuracy;
+  • 4   response is largely relevant and accurate;
+  • 5   response is a perfect match to the actual answer.
+We assess the relationship between human evaluation scores and other metrics using both Pearson and Kendall’s Tau correlation coefficients (see Figure 8 and Table 13).
+Pearson correlation captures linear relationships, while Kendall’s Tau assesses rank-based similarity, making them complementary for evaluating metric alignment.
+Across both measures, F1 Score and LLM-as-a-Judge consistently show the highest positive correlations with average human ratings—indicating not only strong agreement in scale but also in ranking.
+This indicates that both F1 Score and LLM-as-a-Judge closely align with human evaluations, supporting their use as strong performance metrics.
+Metric prioritization and interpretation
+Based on the observed correlations with human judgments, we recommend interpreting LLM-as-a-Judge as the primary indicator of semantic correctness, particularly for higher-complexity questions where answers may be paraphrased, partially implicit, or expressed with variable surface forms.
+F1 score should be viewed as a complementary metric that captures extractive fidelity and coverage, and is most informative when exact grounding in the document is required.
+Discrepancies between the two metrics are therefore not treated as inconsistencies but as signals of different model behaviors (e.g., cautious but semantically accurate answers versus broader but partially grounded ones), as discussed in Section 3.3.
+For practical use, we recommend reporting both metrics jointly: LLM-as-a-Judge to assess answer correctness from a human-aligned perspective, and F1 to diagnose retrieval completeness and over-/under-generation.
+The remaining metrics (Edit Distance and Cosine Similarity) provide additional diagnostic insight into surface-level mismatch and semantic drift but are not intended to be used in isolation for model ranking.
+Batch QA Prompt [SYSTEM INPUT] You are a financial expert, and your task is to answer the questions given to you in batches about the provided credit agreement.
 If you believe the answer is not present in the agreement, say ’Not found’.
-[EXPECTED OUTPUT]
-Respond ONLY with valid JSON, nothing else. See the example below.
-The given document:
-Tesla, Inc. is an American electric vehicle and clean energy company founded in 2003 by Martin Eberhard and Marc Tarpenning. Elon Musk became the largest investor and later CEO.
-The given questions:
-Q1: Who founded Tesla?
+[EXPECTED OUTPUT] Respond ONLY with valid JSON, nothing else.
+See the example below.
+The given document: Tesla, Inc.
+is an American electric vehicle and clean energy company founded in 2003 by Martin Eberhard and Marc Tarpenning.
+Elon Musk became the largest investor and later CEO.
+The given questions: Q1: Who founded Tesla?
 Q2: What year was Tesla founded?
-The expected output:
-    "answers": [               {"question_index": 1,               "answer": "Martin Eberhard,                                 Marc Tarpenning"},               {"question_index": 2,               "answer": "2003"}        ] }
-[USER INPUT]
-{Document Context}
-[QUESTIONS]
-Q1: {question}
-Q2: {question}
+The expected output: { "answers": [       {"question_index": 1,       "answer": "Martin Eberhard,               Marc Tarpenning"},       {"question_index": 2,       "answer": "2003"}    ] } [USER INPUT] {Document Context} [QUESTIONS] Q1: {question} Q2: {question}
 (a) Example of a batch prompt combining up to 50 questions with a single document chunk. The LLM is instructed to return only JSON-formatted answers, each linked to a question index.
-Merging Prompt [SYSTEM INPUT]
-You are a financial expert, and your task is to combine or merge the provided partial answers, coming from different chunks of a credit agreement, into a single final answer for each of the questions given to you.
+Merging Prompt [SYSTEM INPUT] You are a financial expert, and your task is to combine or merge the provided partial answers, coming from different chunks of a credit agreement, into a single final answer for each of the questions given to you.
 If you believe the answer is not present in the agreement, say ’Not found’.
-[EXPECTED OUTPUT]
-The given document:
-Tesla, Inc. is an American electric vehicle and clean energy company founded in 2003 by Martin Eberhard and Marc Tarpenning. Elon Musk became the largest investor and later CEO.
-The given questions:
-Q1: Who founded Tesla?
+[EXPECTED OUTPUT] The given document: Tesla, Inc.
+is an American electric vehicle and clean energy company founded in 2003 by Martin Eberhard and Marc Tarpenning.
+Elon Musk became the largest investor and later CEO.
+The given questions: Q1: Who founded Tesla?
 Q2: What year was Tesla founded?
-The expected output:
-    "answers": [               {"question_index": 1,               "answer": "Martin Eberhard,                                 Marc Tarpenning"},               {"question_index": 2,               "answer": "2003"}        ] }
-[USER INPUT]
-Chunk 1 partial answer JSON: {ans 1}
-Chunk 2 partial answer JSON: {ans 2}
-[QUESTIONS]
-Q1: {question}
-Q2: {question}
+The expected output: { "answers": [       {"question_index": 1,       "answer": "Martin Eberhard,               Marc Tarpenning"},       {"question_index": 2,       "answer": "2003"}    ] } [USER INPUT] Chunk 1 partial answer JSON: {ans 1} Chunk 2 partial answer JSON: {ans 2} [QUESTIONS] Q1: {question} Q2: {question}
 (b) Example of a merging prompt that consolidates partial answers from multiple document chunks into one final JSON-formatted answer per question.
@@
 For multi-chunk documents, a merging prompt consolidates partial answers.
-We implement a two-stage prompting pipeline for QA inference over long documents.
-In the first stage, the document is either processed in full or split into chunks (depending on the context limit of the LLM), and each chunk is paired with a batch of up to 50 questions.
-We then apply a one-shot, instruction-based prompt (Figure 7(a)) that instructs the LLM to act as a financial expert and return an array of answers – either grounded in the document or explicitly stating “Not found" – in valid JSON format.
-The expected format is enforced both through the prompt instructions and through structured parsing logic.
-The parser first attempts to extract a JSON object from the model’s response, supporting both raw JSON and code-fenced output.
-If parsing fails, the system falls back to extracting answers from list-formatted outputs or reattempts the query, up to a maximum number of retries.
-All models are queried with temperature=0.0 for determinism.
-In the second stage, for questions where answers may be distributed across multiple chunks, a merging prompt is issued (Figure 7(b)).
-This prompt consolidates partial answers from earlier responses and asks the model to output a single final answer per question, again using the same strict JSON format.
-This strategy allows the model to incorporate information across multiple views of the document.
-Regarding resources, as mentioned in 4.1, for open-source models we are using Together.AI for some of the models. For the other open-source models, we used 1 H200 GPU.
-D.2 Evaluation Metrics
-We employ four complementary metrics to evaluate model responses.
-  • Word-Level F1 Score  Token-level precision and recall against the gold answer, ensures strict factual correctness in extractive question answering.
-  • Normalized Levenshtein Distance  Character-level string distance, indicating surface-level divergence (described by Yujian and Bo (2007)).
-  • Cosine Similarity  Evaluates the semantic alignment, providing robustness to minor changes.
-  • LLM-as-a-Judge  Gemini-2.0-Pro is prompted to score each prediction from 1 to 5 based on its semantic correctness, given the question and reference answer.
-We exclude ROUGE, BLEU, and similar metrics because they are known to perform poorly on short factual answers, tend to over-penalize surface-level variations and are not well-aligned with human judgment in our setting (Blagec et al., 2022; Chaganty et al., 2018).
 Table 12: Comparison of model performance across difficulty levels, including F1 Score, Normalized Edit Distance, Cosine Similarity, and LLM-as-a-Judge.
@@
 Open-source LLMs
+DeepSeek-R1 0.6269 0.3510 0.6208 3.6965 0.3293 0.5191 0.3179 2.6354 0.2609 0.7733 0.1451 2.0736
 DeepSeek-V3 0.6218 0.3722 0.6077 3.5668 0.4796 0.4912 0.4678 3.0210 0.0856 0.7975 0.0737 1.4795
+GPT OSS 120B 0.5885 0.4187 0.5776 3.5839 0.4684 0.4765 0.4740 3.1661 0.1136 0.7067 0.1202 1.4888
 Llama-4-Maverick-Instruct 0.5552 0.3585 0.5494 3.4482 0.3848 0.5606 0.3669 2.6976 0.2031 0.7231 0.1172 3.1351
@@
 Llama-3.1-Instruct 0.3891 0.5617 0.3793 2.6996 0.2962 0.6255 0.2856 2.2332 0.1169 0.7980 0.0915 1.4315
+Qwen3-Instruct 0.5243 0.4243 0.5133 3.2693 0.2760 0.5795 0.2670 2.4460 0.1240 0.7473 0.1129 1.9578
 Qwen2-Instruct 0.3886 0.6369 0.3710 3.3884 0.3100 0.7031 0.2986 2.9078 0.0762 0.8200 0.0690 1.3288
-D.3 Comparison with Oracle and RAG
-We compare three retrieval settings for answering questions from credit agreements: (1) Oracle, which uses the pieces of text containing the answer; (2) RAG (Retrieval-Augmented Generation), which selects top passages using a retriever; and (3) Full, where the entire agreement is provided to the model.
-In table 13, we observe that RAG performs drastically worse than the other two settings across all difficulty levels, which highlights the incompatibility of standard retrieval techniques with dense, cross-referenced financial documents, often requiring the aggregation of logically related but spatially distant elements.
-Interestingly, the comparison between Oracle and Full Document settings reveals a non-trivial pattern.
-For easy look-up questions, the Oracle setting performs better, which aligns with expectations since such questions benefit from isolating the minimal, most relevant snippet.
-However, for medium and hard questions, the Full Document consistently outperforms the Oracle setting.
-This suggests that isolating the “correct" snippet removes necessary contextual cues that models need for logical, multi-hop inference.
-These findings reinforce the value of credit agreements as a testbed for long-context LLMs, and demonstrate how our framework’s question complexity help uncover the precise scenarios where models struggle, even when given the exact ground-truth passage (see Section 5 and E for the detailed analysis).
-Table 13: Performance comparison across three retrieval settings – Oracle (gold snippet only), RAG (retriever-selected passages), and Full (entire credit agreement) – for each question complexity group.
+(a) Pearson correlation between various metrics. F1 Score and Avg LLM-Human Rating (r = 0.941), and LLM-as-a-Judge and Avg LLM-Human Rating (r = 0.947) show the highest correlations in the matrix.
+(b) Kendall-Tau correlation between various metrics.
+The highest correlations are observed between F1 Score and Avg LLM-Human Rating ($\tau$ = 0.870), and between LLM-as-a-Judge and Avg LLM-Human Rating ($\tau$ = 0.867).
+Figure 8: Pearson and Kendall’s Tau Correlation Matrices.
+Table 13: Pearson and Kendall’s Tau Correlation Tables.
+Pearson Correlation Matrix
+F1 Score Cosine Similarity Edit Distance Sim LLM-as-a-Judge Avg LLM-Human Rating
+F1 Score 1.000 0.933 -0.894 0.922 0.941
+Cosine Similarity 0.933 1.000 -0.910 0.935 0.922
+Edit Distance Sim -0.894 -0.910 1.000 -0.844 -0.862
+LLM-as-a-Judge 0.922 0.935 -0.844 1.000 0.947
+Avg LLM-Human Rating 0.941 0.922 -0.862 0.947 1.000
+Kendall’s Tau Correlation Matrix
+F1 Score Cosine Similarity Edit Distance Sim LLM-as-a-Judge Avg LLM-Human Rating
+F1 Score 1.000 0.902 -0.745 0.854 0.870
+Cosine Similarity 0.902 1.000 -0.746 0.894 0.858
+Edit Distance Sim -0.745 -0.746 1.000 -0.734 -0.733
+LLM-as-a-Judge 0.854 0.894 -0.734 1.000 0.867
+Avg LLM-Human Rating 0.870 0.858 -0.733 0.867 1.000
+D.4 Cost-Constrained Models
+For the cost-constrained models, we used stratified sampling over our question templates.
+Because each template corresponds to a unique point in the (H, P, #SO) space, sampling proportionally over templates preserves the difficulty distribution, the multi-hop vs single-hop ratio, and the relative frequency of set-operation queries.
+Using DeepSeek-V3, the strongest open-weight model evaluated on the full dataset, we have verified that performance on the sampled subset correlates strongly with full-set performance: 0.87 for F-score, 0.84 for edit distance, and 0.88 for cosine similarity.
+D.5 Chunk-Based Evaluation
+As seen in Table 14, the effect of chunking becomes noticeable as the question complexity grows.
+The performance increase is driven mainly by a reduction in abstentions, with models more often returning explicit answers instead of ‘‘Not Found’’.
+However, even for medium questions nearly half is still unanswerable without chunking: it does not reflect improved reasoning or extraction quality, which means chunking does not affect our error analysis or the conclusions drawn in the paper.
+Table 14: Evaluation of Gemini-2.0-Flash with chunking (as in the main paper) and without chunking.
+The context length of this model is large enough to incorporate an entire credit agreement in one prompt.
+Strategy F1 $\uparrow$ Edit $\downarrow$ Cos $\uparrow$ LLM-J $\uparrow$
+Easy
+With 0.6748 0.2679 0.6650 4.0000
+Without 0.7105 0.2987 0.6949 3.9269
+Medium
+With 0.3294 0.6280 0.3121 2.4820
+Without 0.5558 0.4051 0.5452 3.4344
+Hard
+With 0.1172 0.7238 0.0608 2.2838
+Without 0.2565 0.6310 0.2504 2.1088
+D.6 Single Question Evaluation
+We note that the absolute metric values in Table 15 are not representative of task complexity, as a sampling of 20 questions per complexity level was required due to the high cost of single-question evaluation for long-context documents.
+Examining relative differences, the single-question evaluation provides marginal gains for easy and medium questions, but not for hard ones.
+These variations are minor and leave the overall performance patterns across complexity levels unchanged.
+Table 15: Evaluation of Gemini-2.0-Flash in batches of 50 questions (as in the main paper) and with a simple questions.
+Difficulty F1 $\uparrow$ Edit $\downarrow$ Cos $\uparrow$ LLM-J $\uparrow$
+Easy
+Batch 0.5277 0.3773 0.5249 3.7500
+Single 0.5719 0.3573 0.5688 3.8000
+Medium
+Batch 0.3980 0.4220 0.3958 3.2040
+Single 0.4451 0.3695 0.4427 3.5510
+Hard
+Batch 0.5333 0.4223 0.5333 3.1333
+Single 0.4971 0.4367 0.4938 3.0666
+D.7 Human Performance Evaluation
+To establish a reference point for model comparison, we conducted a human evaluation on a balanced subset of 100 questions (20 per difficulty level).
+Three independent human evaluators participated in this study: one annotator who was fully familiar with the documents, and two participants with general financial knowledge but no prior exposure to the documents.
+Their responses were normalized and compared against ground truth using three similarity metrics: word level F1, edit distance and cosine similarity.
+Unlike model performance, which typically declines as task complexity increases, human performance remained relatively stable across levels, as seen in Table 16.
+This suggests that humans effectively leverage contextual reasoning and background knowledge to maintain accuracy even in multi-hop or compositional settings, capabilities that current models still struggle with.
+Table 16: Mean human performance across question levels.
+Level F1 Edit Dist. Cosine Sim.
+1 0.846 0.142 0.834
+2 0.778 0.264 0.750
+3 0.780 0.235 0.756
+4 0.803 0.244 0.795
+5 0.935 0.101 0.928
+Table 17: Performance comparison across four retrieval settings – Full (entire credit agreement), Oracle (gold snippet only), RAG and dynamic RAG (retriever-selected passages) – for each question complexity group.
 Easy Medium Hard
@@
 $\uparrow$ a-Judge $\uparrow$
-Full 0.8542 0.1252 0.8507 4.4893 0.6936 0.3417 0.6932 3.8983 0.1172 0.7238 0.0608 2.2838
-Oracle 0.9113 0.0673 0.9076 4.6950 0.6335 0.4075 0.6360 4.0508 0.1128 0.8067 0.0736 1.9189
-RAG 0.1481 0.7341 0.1473 1.6099 0.1277 0.7543 0.1251 1.5254 0.0191 0.8450 0.0107 1.2027
-D.4 Human Evaluation Metrics
-We conducted a human evaluation of the LLM-generated responses using three independent evaluators.
-Two of these evaluators were annotators familiar with credit agreements, while the third evaluator had no prior exposure to the documents.
-Each evaluator was requested to rate the LLM responses on a scale from 1 to 5, according to the criteria below.
-A total of 100 questions were sampled from multiple templates to ensure maximal diversity in the human evaluation task.
-  • 1   response is irrelevant or does not match the actual answer;
-  • 2   response is somewhat relevant but does not match the actual answer;
-  • 3   response shows slight similarity but lacks accuracy;
-  • 4   response is largely relevant and accurate;
-  • 5   response is a perfect match to the actual answer.
-We assess the relationship between human evaluation scores and other metrics using both Pearson and Kendall’s Tau correlation coefficients (see Figure 8 and Table 14).
-Pearson correlation captures linear relationships, while Kendall’s Tau assesses rank-based similarity, making them complementary for evaluating metric alignment.
-Across both measures, F1 Score and LLM-as-a-Judge consistently show the highest positive correlations with average human ratings—indicating not only strong agreement in scale but also in ranking.
-This indicates that both F1 Score and LLM-as-a-Judge closely align with human evaluations, supporting their use as strong performance metrics.
-Table 14: Kendall’s Tau and Pearson Correlation Tables.
-Kendall’s Tau Correlation Matrix
-F1 Score Cosine Similarity Edit Distance Sim LLM-as-a-Judge Avg LLM-Human Rating
-F1 Score 1.000 0.902 -0.745 0.854 0.870
-Cosine Similarity 0.902 1.000 -0.746 0.894 0.858
-Edit Distance Sim -0.745 -0.746 1.000 -0.734 -0.733
-LLM-as-a-Judge 0.854 0.894 -0.734 1.000 0.867
-Avg LLM-Human Rating 0.870 0.858 -0.733 0.867 1.000
-Pearson Correlation Matrix
-F1 Score Cosine Similarity Edit Distance Sim LLM-as-a-Judge Avg LLM-Human Rating
-F1 Score 1.000 0.933 -0.894 0.922 0.941
-Cosine Similarity 0.933 1.000 -0.910 0.935 0.922
-Edit Distance Sim -0.894 -0.910 1.000 -0.844 -0.862
-LLM-as-a-Judge 0.922 0.935 -0.844 1.000 0.947
-Avg LLM-Human Rating 0.941 0.922 -0.862 0.947 1.000
-(a) Pearson correlation between various metrics. F1 Score and Avg LLM-Human Rating (r = 0.941), and LLM-as-a-Judge and Avg LLM-Human Rating (r = 0.947) show the highest correlations in the matrix.
-(b) Kendall-Tau correlation between various metrics.
-The highest correlations are observed between F1 Score and Avg LLM-Human Rating ($\tau$ = 0.870), and between LLM-as-a-Judge and Avg LLM-Human Rating ($\tau$ = 0.867).
-Figure 8: Kendall’s Tau and Pearson Correlation Matrices.
+Full 0.6747 0.2678 0.6650 4.000 0.3294 0.6279 0.3121 2.4820 0.1172 0.7237 0.0607 2.2837
+Oracle 0.7134 0.2588 0.7021 3.9862 0.3294 0.5986 0.3201 2.6497 0.1127 0.8066 0.0735 1.9189
+RAG 0.1579 0.7364 0.1546 1.6931 0.0695 0.8070 0.0651 1.4131 0.0190 0.8449 0.0107 1.2027
+D-RAG 0.1671 0.7304 0.1631 1.7185 0.1118 0.7693 0.1087 1.5566 0.0107 0.8431 0.0069 1.1088
+D.8 Comparison with Oracle and RAG
+We compare four retrieval settings for answering questions from credit agreements: (1) Full, where the entire agreement is provided to the model, (2) Oracle, which uses the pieces of text containing the answer; (3) RAG (Retrieval-Augmented Generation), which selects top passages using a retriever; and (4) Dynamic RAG, which iteratively decomposes the question into sub-queries, retrieves relevant passages across multiple steps, and aggregates the retrieved evidence before generating the final answer.
+In Table 17, we observe that RAG, both simple and dynamic, performs drastically worse than the other two settings across all difficulty levels, which highlights the incompatibility of standard retrieval techniques with dense, cross-referenced financial documents, often requiring the aggregation of logically related but spatially distant elements.
+Interestingly, the comparison between Oracle and Full Document settings reveals a non-trivial pattern.
+For easy look-up and for medium questions, the Oracle setting performs better, as expected, since such questions benefit from isolating the minimal, most relevant snippet.
+However, for hard questions, the Full Document consistently outperforms the Oracle setting.
+This suggests that isolating the “correct” snippet removes necessary contextual cues that models need for reasoning, which is detailed in the Implicit Information Gaps paragraph of our Error Analysis (Section 4.2).
+These findings reinforce the value of credit agreements as a testbed for long-context LLMs, and demonstrate how our framework’s question complexity help uncover the precise scenarios where models struggle, even when given the exact ground-truth passage (see Section 4 and E for the detailed analysis).
 Appendix E Error Analysis
-Below are detailed examples of errors from each of the aforementioned error categories.
+Detailed examples of errors from each of the aforementioned categories are shown in Table 18 and Figure 9. Aggregate trends across error types and complexity are shown in Figures 10 and 11.
 Table 18: Error analysis with examples.
@@
 While City is factually accurate, the question requires identifying the Location Type of "Dubai, UAE" in the context of the credit agreement.
-The expected answer, zonal office, reflects a more precise understanding of this context.
+The expected answer, zonal office, reflects a more precise understanding.
 Question In what organizations does Hamish Sandhu work?
@@
 Error Explanation Implicit Information Gaps.
-The LLM fails to infer implicit information due to structural complexity in the document.
-In cases involving multiple signatories or nested signature panels, human readers can intuitively associate individuals with their respective organizations based on layout and context.
-However, this relationship is not explicitly stated, and the LLM, lacking this inferential reasoning, incorrectly attributes Hamish Sandhu to the wrong organization.
+The LLM fails to infer implicit information due to structural complexity.
+Human readers can associate individuals with their organizations based on layout and context.
+The LLM incorrectly attributes Hamish Sandhu to the wrong organization.
 Question What is the position held by Alexander Wu but not by Chwan Ming Ho or Edward Chen?
@@
 Error Explanation Set Operation Failures.
-The LLM struggles with this query due to the complexity introduced by multiple set operations, in this case, requiring a comparison of roles across three individuals.
-Alexander Wu is the SVP and General Manager, while Chwan-Ming Ho and Edward Chen are both VPs and General Managers (see figure  9).
-While "SVP" is correct, the inclusion of "General Manager" indicates a failure to correctly subtract shared titles.
+The LLM struggles with multiple set operations.
+Alexander Wu is SVP and General Manager, while Chwan-Ming Ho and Edward Chen are both VPs and General Managers.
+The extra "General Manager" indicates a failure in subtracting shared titles.
 Question What is the role of Celanese Chemicals, Inc. in the agreement?
@@
 Error Explanation Long-Context Retrieval Limitations.
-The model fails to retrieve the correct role because the organizational label "subsidiary" is not explicitly stated in a natural sentence.
+The model fails because "subsidiary" is not explicitly stated in a natural sentence.
 Instead, Celanese Chemicals, Inc.
-appears beneath a heading or clause indicating subsidiaries, often following the signature panel.
-This implicit semantic link requires structural and hierarchical reasoning that current LLMs struggle with—particularly when role information is implied by layout rather than expressed textually.
+appears under headings or clauses indicating subsidiaries, requiring structural and hierarchical reasoning.
 (a) Multiple entities and roles are listed in close proximity, making it difficult for the model to associate roles accurately.
@@
 With increase in number of hops or set operations, the percentage distribution of each error type increases.
-Appendix F Dataset Release and Leaderboard Setup
-Table 16: Dataset statistics of the constructed KG-QAGen-D dataset.
+Appendix F Ablation Study on Medical Documents
+To illustrate the generalizability and applicability of our framework to other domain, we conducted a minor ablation study on the publicly available [^8] https://clinicalinfo.hiv.gov/en/guidelines/hiv-clinical-guidelines-adult-and-adolescent-opportunistic-infections/whats-new "Guidelines for the Prevention and Treatment of Opportunistic Infections in Adults and Adolescents With HIV".
+We extracted the introductory part, the first 10 descriptions of diseases, and the conclusive part, which results in a 225-page document.
+We annotated the document according to the schema in Figure 12, which allowed us to extract questions of various complexity levels over the same 3 dimensions as in the main study (see Tables 20-22 for the full list of question templates).
+Figure 12: Annotation schema of our ablation study.
+We ran our experiments with the best-performing open-weight model, DeepSeek-V3.
+In Table 19, we observe a clear performance decrease as the question complexity increases, which illustrates that our framework is generalizable and applicable to other domains.
+Table 19: Evaluation of DeepSeek-V3 on the ablation study dataset.
+Difficulty F1 $\uparrow$ Edit $\downarrow$ Cos $\uparrow$ LLM-J $\uparrow$
+Easy 0.2585 0.6169 0.2522 2.2249
+Medium 0.0753 0.7512 0.0662 1.5587
+Hard 0.0 0.8417 0.0 1.0681
+Table 20: Ablation 1-hop Question Templates
+Template P H #SO
+What pathogen is transmitted by [Vector]? [if one] 0 1 0
+What vector transmits [Pathogen]? [if one] 0 1 0
+What pathogen causes [Disease]? [if one] 0 1 0
+What disease is caused by [Pathogen]? [if one] 0 1 0
+What disease is treated by [Medication]? [if one] 0 1 0
+What medication treats [Disease]? [if one] 0 1 0
+What pathogens are transmitted by [Vector]? 1 1 0
+What vectors transmit [Pathogen]? 1 1 0
+What pathogens cause [Disease]? 1 1 0
+What diseases are caused by [Pathogen]? 1 1 0
+What diseases are treated by [Medication]? 1 1 0
+What medications treat [Disease]? 1 1 0
+What pathogen is transmitted by [Vector 1] and [Vector 2]? [if one] 0 1 1
+What vector transmits [Pathogen 1] and [Pathogen 2]? [if one] 0 1 1
+What pathogen causes [Disease 1] and [Disease 2]? [if one] 0 1 1
+What disease is caused by [Pathogen 1] and [Pathogen 2]? [if one] 0 1 1
+What disease is treated by [Medication 1] and [Medication 2]? [if one] 0 1 1
+What medication treats [Disease 1] and [Disease 2]? [if one] 0 1 1
+What pathogens are transmitted by [Vector 1] and [Vector 2]? 1 1 1
+What vectors transmit [Pathogen 1] and [Pathogen 2]? 1 1 1
+What pathogens cause [Disease 1] and [Disease 2]? 1 1 1
+What diseases are caused by [Pathogen 1] and [Pathogen 2]? 1 1 1
+What diseases are treated by [Medication 1] and [Medication 2]? 1 1 1
+What medications treat [Disease 1] and [Disease 2]? 1 1 1
+What pathogen is transmitted by [Vector 1] but not [Vector 2]? [if one] 0 1 2
+What vector transmits [Pathogen 1] but not [Pathogen 2]? [if one] 0 1 2
+What pathogen causes [Disease 1] but not [Disease 2]? [if one] 0 1 2
+What disease is caused by [Pathogen 1] but not [Pathogen 2]? [if one] 0 1 2
+What disease is treated by [Medication 1] but not [Medication 2]? [if one] 0 1 2
+What medication treats [Disease 1] but not [Disease 2]? [if one] 0 1 2
+What pathogens are transmitted by [Vector 1] but not [Vector 2]? 1 1 2
+What vectors transmit [Pathogen 1] but not [Pathogen 2]? 1 1 2
+What pathogens cause [Disease 1] but not [Disease 2]? 1 1 2
+What diseases are caused by [Pathogen 1] but not [Pathogen 2]? 1 1 2
+What diseases are treated by [Medication 1] but not [Medication 2]? 1 1 2
+What medications treat [Disease 1] but not [Disease 2]? 1 1 2
+Table 21: Ablation 2-hop Question Templates
+Template P H #SO
+What vector transmits a pathogen which causes [Disease]? [if one] 0 2 0
+What disease is caused by a pathogen which is transmitted by [Vector]? [if one] 0 2 0
+What medication treats a disease which is caused by [Pathogen]? [if one] 0 2 0
+What pathogen causes a disease which is treated by [Medication]? [if one] 0 2 0
+What vectors transmit a pathogen which causes [Disease]? 1 2 0
+What diseases are caused by a pathogen which is transmitted by [Vector]? 1 2 0
+What medications treat a disease which is caused by [Pathogen]? 1 2 0
+What pathogens cause a disease which is treated by [Medication]? 1 2 0
+What vector transmits a pathogen which causes [Disease 1] and [Disease 2]? [if one] 0 2 1
+What disease is caused by a pathogen which is transmitted by [Vector 1] and [Vector 2]? [if one] 0 2 1
+What medication treats a disease which is caused by [Pathogen 1] and [Pathogen 2]? [if one] 0 2 1
+What pathogen causes a disease which is treated by [Medication 1] and [Medication 2]? [if one] 0 2 1
+What vectors transmit a pathogen which causes [Disease 1] and [Disease 2]? 1 2 1
+What diseases are caused by a pathogen which is transmitted by [Vector 1] and [Vector 2]? 1 2 1
+What medications treat a disease which is caused by [Pathogen 1] and [Pathogen 2]? 1 2 1
+What pathogens cause a disease which is treated by [Medication 1] and [Medication 2]? 1 2 1
+What vector transmits a pathogen which causes [Disease 1] but not [Disease 2]? [if one] 0 2 2
+What disease is caused by a pathogen which is transmitted by [Vector 1] but not [Vector 2]? [if one] 0 2 2
+What medication treats a disease which is caused by [Pathogen 1] but not [Pathogen 2]? [if one] 0 2 2
+What pathogen causes a disease which is treated by [Medication 1] but not [Medication 2]? [if one] 0 2 2
+What vectors transmit a pathogen which causes [Disease 1] but not [Disease 2]? 1 2 2
+What diseases are caused by a pathogen which is transmitted by [Vector 1] but not [Vector 2]? 1 2 2
+What medications treat a disease which is caused by [Pathogen 1] but not [Pathogen 2]? 1 2 2
+What pathogens cause a disease which is treated by [Medication 1] but not [Medication 2]? 1 2 2
+Table 22: Ablation 3-hop Question Templates
+Template P H #SO
+What vector transmits a pathogen which causes a disease which is treated by [Medication]? [if one] 0 3 0
+What medication treats a disease which is caused by a pathogen which is transmitted by [Vector]? [if one] 0 3 0
+What vectors transmit a pathogen which causes a disease which is treated by [Medication]? 1 3 0
+What medications treat a disease which is caused by a pathogen which is transmitted by [Vector]? 1 3 0
+What vector transmits a pathogen which causes a disease which is treated by [Medication 1] and [Medication 2]? [if one] 0 3 1
+What medication treats a disease which is caused by a pathogen which is transmitted by [Vector 1] and [Vector 2]? [if one] 0 3 1
+What vectors transmit a pathogen which causes a disease which is treated by [Medication 1] and [Medication 2]? 1 3 1
+What medications treat a disease which is caused by a pathogen which is transmitted by [Vector 1] and [Vector 2]? 1 3 1
+What vector transmits a pathogen which causes a disease which is treated by [Medication 1] but not [Medication 2]? [if one] 0 3 2
+What medication treats a disease which is caused by a pathogen which is transmitted by [Vector 1] but not [Vector 2]? [if one] 0 3 2
+What vectors transmit a pathogen which causes a disease which is treated by [Medication 1] but not [Medication 2]? 1 3 2
+What medications treat a disease which is caused by a pathogen which is transmitted by [Vector 1] but not [Vector 2]? 1 3 2
+Appendix G Dataset Release and Leaderboard Setup
+To facilitate reproducibility and future research (Tatarinov et al., 2025), we are publicly releasing (under CC-BY NC ND 4.0 license) QA pairs for 40 documents out of the total 170 as the development set, allowing researchers to validate model performance under controlled conditions.
+The remaining QA pairs serve as the test set, which is not publicly available to prevent data contamination (Sainz et al., 2023).
+For this test set, we are making the questions public, without the ground-truth answers.
+The distribution of the test and dev sets is provided in Table 23.
+In addition, we are hosting a leaderboard [^9] https://huggingface.co/spaces/gtfintechlab/KG-MuLQA-D-Leaderboard, following standard practices in recent LLM evaluation benchmarks (Galarnyk et al., 2025; Yue et al., 2024; Zhao et al., 2024a; Lu et al., 2023).
+Table 23: Dataset statistics of the constructed KG-MuLQA-D dataset.
 Stats Dev Test Total
 # Documents 40 130 170
-[…]
+[2pt/2pt] # Question per doc (min) 1 1 1
 # Question per doc (avg) 14.75 23.49 21.44
 # Question per doc (max) 83 428 428
-[…]
+[2pt/2pt] # Easy Questions (count) 1,499 5,051 6,550
 # Medium Questions (count) 2,680 10,203 12,883
 # Hard Questions (count) 239 467 706
-[…]
-To facilitate reproducibility and future research, we are publicly releasing (under CC-BY NC ND 4.0 license) QA pairs for 40 documents out of the total 170 as the development set, allowing researchers to validate model performance under controlled conditions.
-The remaining QA pairs serve as the test set, which is not publicly available to prevent data contamination (Sainz et al., 2023).
-Instead, we (1) are making the test set questions public, without the ground-truth answers, and (2) will host an online evaluation platform on HuggingFace [^6] https://huggingface.co/, enabling researchers to submit model predictions for the test set and to benchmark performance through a leaderboard.
-This setup follows standard practices in recent LLM evaluation benchmarks (Lu et al., 2023; Yue et al., 2024; Zhao et al., 2024a), ensuring that models are assessed under zero-shot conditions without access to test set ground-truth answers.
-The distribution of the test and dev samples is provided in Table 16.
-Appendix G Ethics Statement
-We generally avoid mentioning specific agreements and their details.
-Instead, we use placeholders such as ABCDE Inc.
-for organization names, Richard Roe for person names, and 10 Peachtree...
-for locations to illustrate the structure of agreement entities and their relations.
-However, in our annotation guide (Appendix A) and error analysis (Appendix E), we include screenshots of real document annotations and mention specific examples to highlight the complexity of these agreements.
-Moreover, we store the raw HTML files of the agreements in our GitHub repository, enabling other researchers to reproduce our benchmark or conduct further experiments.
-As these documents are publicly available through the SEC’s EDGAR system, their inclusion for both illustrative and experimental purposes raises no ethical concerns.
-Appendix H Acknowledgments
-We would like to thank Archishman V B, Jorge Navarro, Krish Shah, Kritika Bansal, Meher Bhardwaj, Nilesh G, Rushi Glasswala, Soumyajit Basu, Tvisha Shah and Xu Huang for contributing to the initial round of annotations and early-stage coding experiments.
-Appendix I Author Contributions
-NT led the data acquisition by extracting credit agreements from SEC EDGAR, developed the annotation guidelines specifying the entity and relation schemas, conceptualized the RDF-based knowledge graph framework, devised multi-dimensional QA templates, and benchmarked long-context LLMs in both RAG and oracle settings.
+[2pt/2pt] # Questions (count) 4,418 15,721 20,139
+Appendix H Author Contribution
+NT led the data acquisition by extracting credit agreements from SEC EDGAR, developed the annotation guidelines specifying the entity and relation schemas, conceptualized the RDF-based knowledge graph framework, devised multi‑dimensional QA templates, and benchmarked long‑context LLMs in both RAG and oracle settings.
 VK extensively annotated and cleaned documents, identified edge cases and was involved in developing the KG construction and question-answer generation modules.
@@
 HS contributed to the document annotation process, developed the pipelines for KG construction and SPARQL QA generation and assisted in LLM response analysis.
-AR implemented the initial end-to-end benchmarking pipeline and supported the human evaluation.
+AR implemented the initial end‑to‑end benchmarking pipeline and supported the human evaluation.
 He experimented with various prompting techniques and evaluation metrics and created one-shot prompts used in experiments.
-HSA contributed to all annotation stages and SPARQL-based question generation.
+HSA contributed to all annotation stages and SPARQL‑based question generation.
 VS, AL, and RL contributed to annotation, data cleaning, and manuscript refinement.
-AS supervised all iterations of the annotation process, guided edge-case incorporation, shaped research direction for document-level question answering, advised on ablation design, and contributed to error analysis and writing.
+AS supervised all iterations of the annotation process, guided edge‑case incorporation, shaped research direction for document-level question answering, advised on ablation design, and contributed to error analysis and writing.
 SC provided overall project supervision and guidance.
-NeurIPS Paper Checklist
-The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact.
-Do not remove the checklist: The papers not including the checklist will be desk rejected.
-The checklist should follow the references and follow the (optional) supplemental material.
-The checklist does NOT count towards the page limit.
-Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
-  • You should answer [Yes] , [No] , or [N/A] .
-  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
-  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
-The checklist answers are an integral part of your paper submission.
-They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers.
-You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
-The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation.
-While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used").
-In general, answering "[No] " or "[N/A] " is not grounds for rejection.
-While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate.
-All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix.
-If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
-IMPORTANT, please:
-  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
-  • Keep the checklist subsection headings, questions/answers and guidelines below.
-  • Do not modify the questions and only use the provided macros for your answers.
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: The last paragraph of introduction explicitly describes the contributions. The abstract conveys the same idea slightly differently but without overclaiming or contradicting.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: See paragraph “Opportunities and Extensions" in the Discussion & Conclusion section.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory assumptions and proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [N/A]
-  14. Justification: No theoretical layouts, proofs, etc. are contributed.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental result reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19.
-Justification: We open-source the benchmarking codes and a part of the dataset, and mention it in the introduction, in the Dataset Release and Leaderboard Setup appendix.
-In addition, README file in our repositories provide explicit instructions to running the codes and getting the results.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24.
-Justification: We open-source only a part of the data, but we provide links to the data and the benchmarking codes in the header with instructions in the repositories.
-In additional, we provide reasoning to open-sourcing only a part of the data (see Appendix F Dataset Release and Leaderboard Setup).
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental setting/details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: See Appendix D.1 Prompting Strategy for the technical details of implementation, and Appendix F Dataset Release and Leaderboard Setup for the test-dev split statistics.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment statistical significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [N/A]
-  34.
-Justification: The paper focuses on a evaluation of QA pairs, and treats it as deterministic rather than probabilistic, so error estimation is not discussed.
-Having said that, we conduct human evaluation of LLMs’ responses (on a sample) and calculate correlations with the other metrics to see which ones align with the human preferences better.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments compute resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: See the last paragraph of appendix D.1 Prompting Strategy.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code of ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: We believe our work fully conforms with the NeurIPS Code of Ethics. We also provide Ethics Statement (Appendix G) for clarification of using credit agreements.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [N/A]
-  49. Justification: Our work does not introduce new model capabilities or deployment scenarios, so societal impacts are not directly applicable to the scope of this study.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54.
-Justification: We believe our benchmark does not have high risk of misuse since the source data is publicly available documents with no images, personal information, etc.
-We mention that we believe there are no ethics concerns in Appendix G Ethics Statement.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: Everyone who contributed to the project till its final stage became a co-author. The other contributers are acknowledge in Appendix H Acknowledgments.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: Both repositories (for data and codes) have README files with instructions to replicate the benchmarking pipeline.
-  64. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  65. 14.
-    Crowdsourcing and research with human subjects
-  66.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  67. Answer: [N/A]
-  68. Justification: No crowdsourcing or research with human subjects was conducted.
-  69. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  70. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  71.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  72. Answer: [N/A]
-  73. Justification: No crowdsourcing or research with human subjects was conducted.
-  74. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
-  75. 16.
-    Declaration of LLM usage
-  76.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-  77. Answer: [N/A]
-  78. Justification: LLMs were used only for refining the writing, not for a original contribution.
-  79. Guidelines:
-    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
-    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
+Appendix I Acknowledgments
+We would like to thank Anvita Mahajan, Archishman VB, Arvind K R, Daksh Jain, Deng Li, Jorge Navarro Gracia, Kaustubh Gayadhankar, Krish Shah, Kritika Bansal, Meher Bhardwaj, Nilesh Gupta, Piyush Mohapatra, Pranav Krishna, Rushi Glasswala, Shashwat Bajpai, Soumyajit Basu, Tvisha Shah, Vishnu Varma, Xu Huang and Udish Jangid for contributing to the initial round of annotations and early-stage coding experiments.